### v1

# Analyzing Classroom Interaction Data Using Prompt Engineering and Network Analysis

Abstract
Classroom interactions play a vital role in developing critical thinking, collaborative problem-solving abilities, and enhanced learning outcomes. While analyzing these interactions is crucial for improving educational practices, the examination of classroom dialogues presents significant challenges due to the complexity and high-dimensionality of conversational data. This study presents an integrated framework that combines prompt engineering with network analysis to investigate classroom interactions comprehensively. Our approach automates utterance classification through prompt engineering, enabling efficient and scalable dialogue analysis without requiring pre-labeled datasets. The classified interactions are subsequently transformed into network representations, facilitating the analysis of classroom dynamics as structured social networks. To uncover complex interaction patterns and how underlying interaction structures relate to student learning, we utilize network mediation analysis. In this approach, latent interaction structures, derived from the additive and multiplicative effects network (AMEN) model that places students within a latent social space, act as mediators. In particular, we investigate how the gender gap in mathematics performance may be mediated by students’ classroom interaction structures.
Keywords: Classroom Interaction, Prompt Engineering, Network Analysis, Additive and Multiplicative Effect Network Model, Network Mediation Analysis
1 Introduction
1.1 Background: Classroom Dialogue Analysis
Classroom interactions play a crucial role in the learning process and substantially influence students’ motivation, comprehension, and academic achievement. Through effective student-teacher and student-student interactions, learners develop essential skills in critical thinking, collaborative work, and problem-solving. Analysis of these student interactions reveals underlying behavioral patterns that either facilitate or impede effective learning.
The analysis of classroom dialogues provides a powerful lens for examining student interactions, offering a direct understanding of engagement patterns and learning processes. Sedova et al. (2019) demonstrates a significant correlation between participation in substantive dialogue and improved academic performance. Similarly, Mercer (2010) establishes that student conversations foster collaborative problem-solving abilities and critical thinking skills through collective meaning-making. These empirical findings underscore the dual nature of classroom dialogues: they simultaneously manifest student interactions and serve as the primary medium through which these interactions can be systematically investigated.
Conventional Approaches
While classroom conversations offer valuable insights into learning dynamics, their analysis presents considerable analytical challenges due to their inherent complexity and high dimensionality. These conversations are often characterized by dynamic shifts in topic, speaker turns, and the interplay of various communicative functions, making manual analysis particularly difficult. These also encompass speaker characteristics (e.g., their speaking style, experience), emotional tone (e.g., enthusiasm, uncertainty, affect), and context (e.g., type of task, classroom environment). To address such complexity, researchers often simplify the analysis by categorizing the intent of individual messages or utterances using predetermined coding schemes. For example, researchers may classify utterances according to function, such as whether they are questions, answers, elaborations, or agreements (Webb et al., 2014; Franke et al., 2015). This strategy entails classifying utterances based on semantic, syntactic, and pragmatic features, and using these classifications to gain insights into specific aspects of the interaction.
[…]
These limitations highlight the need for automated analytical strategies to thoroughly examine classroom dynamics. Although several deep learning methods have been proposed for automatic dialogue classification (Liu et al., 2017; Raheja and Tetreault, 2019; Song et al., 2021), they depend on large labeled datasets that are difficult to acquire and may not generalize well across different research objectives. For example, Raheja and Tetreault (2019) proposed a context-aware self-attention model for general dialogue act classification, but it requires extensive manual annotation for adaptation to specific tasks. On the other hand, Song et al. (2021) proposed an automated model for classroom dialogue classification by combining a convolution neural network (CNN) and a bidirectional-long short-term memory (Bi-LSTM), but it also requires large labeled datasets and may have difficulty generalizing to diverse research objectives. These challenges motivate the development of alternative approaches that minimize reliance on pre-labeled datasets and offer greater flexibility.
Prompt Engineering
[…]
Strategic prompt engineering enables researchers to classify dialogue interaction types even without traditional training data for fine-tuning, provided the prompts incorporate sufficient contextual information. Nevertheless, while prompt engineering facilitates classification of individual utterances, it cannot fully capture the complex dynamics of classroom interactions. The classification of discrete utterances reveals speaker identities and discussion content but fails to illuminate critical patterns such as interaction frequency, structural relationships, and the relationships between students.
Network Analysis
To address these gaps, network analysis offers a powerful approach for analyzing the structure of student interactions (Grunspan et al., 2014; Bokhove, 2018). By representing the classroom as a network, with students as nodes and their interactions as edges, researchers can analyze complex patterns of communication frequency and type within broader classroom dynamics. For example, centrality metrics can reveal key insights about student influence and prominence within the network (Williams et al., 2019). To uncover deeper patterns of local interactions and latent social dynamics, latent space modeling techniques can be used (Hoff et al., 2002; Hoff, 2005; Sosa and Buitrago, 2021). The projection of students into a latent social space facilitates visualization and analysis of underlying classroom social structures, revealing both student clusters with similar interaction patterns and the influence of these patterns on classroom dynamics and learning outcomes. Network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022) can be used to further investigate how classroom network structures influence student learning outcomes. By using latent vectors from latent space modeling, this approach quantifies the mechanisms through which classroom interactions shape achievement, offering insights into how social structures mediate educational outcomes.
1.2 Purpose and Structure of the Paper
In this paper, we present an integrated framework for analyzing classroom interactions by combining the strengths of prompt engineering and network analysis. Our methodology begins with automated dialogue classification through prompt engineering, addressing the limitations of traditional manual labeling approaches. We employ carefully constructed prompts to guide the generative model in classifying interactions according to predefined criteria, establishing the foundation for network-based analysis of classroom dynamics. We then transform these classified interactions into network representations and analyze them using a network mediation model. Specifically, we employ the additive and multiplicative effects network (AMEN) model (Hoff, 2021) model to extract latent social structures and utilize them as mediators to investigate how classroom networks mediate the relationship between gender and mathematics performance. Through examination of how these latent social structures influence student outcomes, with particular attention to gender differences in mathematics performance, we illuminate the mechanisms by which network structures mediate the impact of classroom dynamics on student achievement. This comprehensive methodological integration can provide educational researchers with a powerful analytical tool for investigating the complex relationships between social dynamics and academic performance.
The remainder of this paper is structured to present our methodology and findings systematically. Section 2 examines our prompt engineering approach, detailing the automated utterance classification process and the creation of labeled interaction data. Section 3 introduces our network analysis framework, beginning with the conceptual foundations of the network mediation model, progressing through the application of the AMEN model for latent position estimation, and concluding our key analytical findings. Section 4 concludes by exploring the broader implications of our research and identifying promising directions for future studies that build upon this integrated methodological framework.
2 Prompt Engineering for Utterance Classification
2.1 Data: Classroom Interaction Dataset
We analyze classroom interaction data from Webb et al. (2023). The dataset includes observations of 20 third-grade students (9 male, 11 female) from a classroom located in a low-income urban community in Southern California. Researchers collected data over a five-month period, during which time they recorded six one-hour math lessons using strategically placed cameras to capture comprehensive classroom interactions between teachers and students, as well as interactions among students. The recordings were carefully transcribed to include contextual information, including speaker identities, timestamps, and job specifications. Initially, students’ mathematics performance was assessed with the California standards test (CST) (EdData, 2013). Due to the statewide transition in the assessment framework, however, the final performance was assessed by the author-developed multiplication and division test (Webb et al., 2023). The pre-test, based on CST, yielded a mean score of 369.35 (SD = 54.31). The post-test, specifically designed to assess problem-solving strategies and accuracy, resulted in a mean score of 15.08 out of 24 (SD = 6.49).
[…]
A significant challenge in analyzing dialogue data arises from the context-dependent nature of utterances. The meaning of conversations often goes beyond individual statements, necessitating a careful examination of the surrounding dialogue. The following brief dialogue example illustrates how seemingly simple responses can convey subtle meanings when considered within their conversational context:
Student A: I think it’s 39 because each one had 39.
Student B: Did you put it one by one?
Student A: Yeah.
Student C: Mine got 35.
[…]
2.2 Prompt Engineering Methodology
The complexity and high-dimensional nature of classroom interactions necessitate a robust method for classifying and analyzing dialogue intent. As discussed in the Introduction Section, existing analytical approaches require manual labeling, which is resource-intensive, while deep learning methods require extensive labeled datasets. Our methodology addresses these issues through prompt engineering, by leveraging pre-trained models to classify dialogue intents accurately without labeled training data or computational fine-tuning. This approach maximizes the capabilities of pre-trained models while maintaining efficiency in resource utilization and processing time.
In-context learning (ICL) is a specialized application of prompt engineering in which pre-trained large language models (LLMs) learn exclusively from information provided within the prompt, eliminating the need for parameter updates. This approach parallels human learning processes through inference while minimizing computational requirements for task adaptation (Dong et al., 2022). Few-shot prompting, one of the ICL strategies, enhances model comprehension by incorporating targeted task examples within the prompt. The effectiveness of few-shot prompts in ICL has been explained through probabilistic frameworks such as hidden Markov models (Baum and Petrie, 1966), where carefully selected examples guide model reasoning by establishing sequential dependencies that reduce prediction uncertainty (Xie et al., 2021). Min et al. (2022) further demonstrates that optimal performance depends not merely on increasing example quantity but on developing prompt formats that precisely align with task-specific requirements, including both structural elements and linguistic formulation.
[…]
Figure 1: System and User Prompts for Utterance Classification. Includes guidelines, category definitions, and example dialogues for analyzing and classifying the contribution of utterances in collaborative problem-solving processes.
Figure 1 illustrates the user and system prompts implemented in this study. The prompts define the utterance classification task, outline categorization rules, and provide illustrative dialogues to clarify distinctions between categories. The user prompt provides clear guidelines for classifying utterances into three main categories: ‘explain own ideas’, ‘engage others’ ideas’, or ‘uncorrelated’. For the ‘engage others’ ideas’ category, it further defines three levels of engagement (low, medium, high) and offers specific criteria for making classification decisions, including special cases and exceptions. The system prompt supports these guidelines by stressing the importance of accurate categorization, emphasizing subtle differences between labels, and ensuring consistency across classifications. To enhance output reliability and reduce variability in the model’s responses, we adjusted the temperature parameter to nearly zero, optimizing consistency in classification.
2.3 Utterance Classification Results
We conducted utterance classification using prompts carefully constructed, as illustrated in Figure 1. These prompts were designed incorporating comprehensive contextual information, including dialogue tasks, categorization rules, and classification guidelines. To optimize classification accuracy, we included five few-shot examples in the prompts, each structured to specify speaker identity, utterance content, classification label, engagement level, and relevant references. Complete prompt specifications, including the few-shot examples, are detailed in the Section C of the Supplementary Material. Instead of individual utterances, dialogue chunks segmented by timestamps from the original dataset were included as the inputs for classification within our prompts. This approach allows the prompts to capture conversational context effectively and inform the classification decisions. The classification process was executed on an NVIDIA RTX A6000 GPU with an Intel(R) Xeon(R) Gold 6426Y CPU in a Linux environment, requiring approximately 1.5 hours for completion.
[…]
  Utterance Type Engagement Level Number of Utterances
Explain Own Idea (EXP) 1,335
Engage Others Idea (EOI) Total 1,251
High Engagement 107
Medium Engagement 553
Low Engagement 591
Uncorrelated 1,544
Total Utterances 4,130
Table 1: Classification Results by Utterance Type and Engagement Level
[…]
  Name texts Label Engagement Level
  Delilah Okay now. First I put 2 days to represent the days. And then I put 70 units in each group (gestures across the two groups on her notebook). explain own idea -
Keila First I… uncorrelated -
Delilah And I came up with 39 because each one had 39. explain own idea -
Keila Did you put it one by one? engage others idea medium
Delilah Yeah. explain own idea -
Keila Mine got 35. engage others idea low
Delilah So I got 1, 2, 3, 4, 5, 6, 7, 8, blah blah blah. (counting the marks in one group) explain own idea -
Keila Did you put 1, 2, 3, 4, 5, 6? (gestures marking in alternating groups as she counts, demonstrating action in solving) engage others idea medium
Delilah Yeah! And 10 divided by 2 is 3. I meant 39. explain own idea -
Table 2: Examples: Identifying labels (explain vs. engage) and engagement levels.
  Name texts Label Engagement Level
  William Okay so first what I did with this is I did 70 and 2. So I had two days. 1 day, 2 days, and I counted all the way up to 70 by 2, and I counted. And so now I got 32. explain own idea -
Julian I disagree with you cause I got 1, 2, 3, (…), 14 (counts his 14 rods) engage others idea high
William Wait hold on. You don’t add 7 plus 7. If you add 7 plus 7 that would be 14. But are you trying to add? Or times? engage others idea high
Table 3: High-Level Engagement in Problem-Solving
Second, Table 3 illustrates the system’s ability to capture complex interaction sequences. The example begins with a student’s explanation of their approach (classified as EXP), followed by a peer’s substantive disagreement (classified as high-level EOI). The subsequent probing question, which builds upon this disagreement, maintains the high-level EOI classification. This sequence demonstrates the capacity of the system to identify sophisticated interactions where students progressively build upon and critically engage with each other’s mathematical reasoning.
  Name texts Label Engagement Level
  Natalie Okay, let me go. uncorrelated -
Kimberly I missed a number. uncorrelated -
Natalie I’m looking at the hundreds chart. (refers to yellow sheet) And then I went to 75. explain own idea -
Kimberly Why did you go to 75? engage others idea medium
Natalie Because that’s how much shells… explain own idea -
Kimberly (talks over N) No, it was 70. engage others idea low
Natalie …that was how much shells she found. explain own idea -
Kimberly It was 70. engage others idea low
Natalie On the second strategy. explain own idea -
Kimberly Oh! uncorrelated -
Natalie And then, I was thinking how could I do it. Then, I was thinking about money, I don’t know why. Then, I remembered a quarter, so then I know what that is. (pause) And then, I started like because I remember the quarters, so I went 20, 25, no… explain own idea -
Kimberly (talking to another group) You should keep on sharing strategies. That’s what we are doing. uncorrelated -
Samantha We are already done. uncorrelated -
Kimberly We are at the strategies, but she and me realized that we strategies on the 200 chart. uncorrelated -
Table 4: Differentiating Relevant and Off-Topic Remarks.
[…]
2.4 Validation of Utterance Classification
For the validation of our prompt engineering approach, we focused on assessing the alignment between intended classifications and actual outcomes. Due to the absence of pre-labeled data in our classroom dialogue dataset, we utilized the DailyDialog dataset (Li et al., 2017) as our external validation source. This dataset provides a structured framework for dialogue act classification, categorizing utterances into four distinct types: inform, questions, directives, and commissive. To evaluate our methodology, we developed dialogue act classification (DAC) specific prompts for this dataset, incorporating five dialogue chunks examples to create few-shot examples prompts. We compared the performance of these few-shot prompts against zero-shot approaches, enabling us to quantify the impact of including dialogue examples in the classification process.
To establish the effectiveness and generalizability of our prompt engineering approach, we conducted comprehensive comparisons with established deep learning models. Context-Aware Self-Attention model (Raheja and Tetreault, 2019, CASA) employs a self-attention mechanism integrated with hierarchical deep neural networks to recognize dialogue acts within conversational contexts. The Deep Co-Interactive Relation Network (Qin et al., 2020, DCR-Net) employs a hierarchical encoder and relation layer to integrate contextual and interaction information for dialogue act classification. A co-interactive graph attention network model (Qin et al., 2021, Co-GAT) jointly performs DAC and sentiment classification, leveraging both contextual and interaction information through cross-utterance and cross-task connections. Additionally, Zhao et al. (2023) presented classification results using prompt engineering with the GPT-3.5-turbo model, and we compared our results against these presented results. We evaluated classification performance using three complementary metrics: macro-F1, weighted-F1 scores, and overall accuracy. Macro-F1 computes the F1 score for each class individually and then averages them, providing a balanced measure of performance across all classes, regardless of their frequencies. This is particularly useful in imbalanced datasets, where certain classes may dominate. On the other hand, Weighted-F1 calculates the F1 score for each class and takes into account the number of true instances for each class, giving more importance to classes with more instances. These metrics help assess the model’s performance from different perspectives, considering both class balance and overall accuracy.
  Model Acc Macro-F1 Weighted-F1
  CASA - 0.78 -
DCR-Net + Co-Attention - 0.79 -
Co-GAT - 0.79 -
ChatGPT 1-shot + P.E 0.67 - 0.65
ChatGPT 3-shot + P.E 0.73 - 0.71
Llama-3.1-8B-Instruct Zero-shot 0.69 0.57 0.68
Llama-3.1-8B-Instruct Few-shot 0.75 0.66 0.74
Table 5: Comparison of Prompt Engineering results with other baselines on the DAC task. Experimental results for CASA, DCR-Net + Co-Attention, and Co-GAT are partly cited from Qin et al. (2021) and Zhao et al. (2023). P.E. stands for prompt engineering where the 1-shot and 3-shot settings use 1 and 3 examples in the prompt, respectively.
Table 5 presents our comparative analysis of prompt engineering approaches and state-of-the-art deep learning models. Results demonstrate that the few-shot prompt engineering method, incorporating five dialogue chunks as examples, achieves substantially improved performance compared to the zero-shot approach. While some comparative studies did not report all evaluation metrics, we have calculated a comprehensive set of metrics for our prompt engineering results to facilitate thorough comparison across different models. This analytical approach ensures equitable performance assessment across varying methodologies.
The comparative analysis reveals that prompt engineering performance, while effective, may not fully match that of deep learning models due to fundamental methodological distinctions. Deep learning approaches benefit from extensive optimization using pre-constructed labeled datasets and generate specialized embeddings tailored to specific tasks. In contrast, prompt engineering operates solely through label descriptions without prior training data, relying on the inherent language understanding capabilities of the model. This distinction becomes particularly significant in contexts where classification criteria exhibit variability due to multiple annotators or crowdsourcing processes. Nevertheless, these differences suggest that prompt engineering may be especially valuable in generating consistent labels, as it helps maintain classification criteria across datasets.
3 Network Analysis of Classroom Interaction
3.1 Network Construction from Classified Utterances
After classifying utterances using prompt engineering, we created a directed network representation to systematically analyze student interaction patterns. The representation of student interaction patterns as networks enables both visualization and systematic analysis of classroom dynamics.
We formalized student interactions using an adjacency matrix, where the matrix elements indicate the frequency of specific interaction types between two students during the observation period. We then visualized the constructed network structure by using nodes to represent individual students and directed weighted edges between nodes to represent the nature and frequency of the interactions between the two students. For instance, when a student offers multiple explanations to a peer, a directed edge is formed from the explaining student (source) to the listening student (target), with the edge weight reflecting the total number of explanatory interactions. Engagement interactions similarly generate directed edges from the engaging student to the idea originator, with edge weights reflecting interaction frequency. To capture the varying intensity of student engagement more precisely, we implemented a differential weighting scheme for these interactions. While explanatory interactions maintained equal weights, we applied multiplication factors to different levels of engagement: medium-level engagements received double weight, and high-level engagements received triple weight relative to low-level engagements. This weighted approach ensures that more substantive forms of engagement exert proportionally greater influence in the network structure, better reflecting their educational significance.
While this network construction process allows us to represent valuable aspects of classroom interactions, analyzing classroom conversation data in a network also presents unique analytical challenges. Unlike conventional social network datasets, our classroom interactions are often characterized by limited and sparse connections, as conversations frequently occur within small groups during specific question discussions. To overcome the sparsity of individual lesson networks and facilitate a more comprehensive analysis of overall interaction patterns, we aggregated data across all six lessons to construct a single composite network.
3.1.1 Network Graphs
Figure 2 illustrates these network transformations, where each node represents a student and edges represent their interactions. Node colors indicate students’ standardized mathematics performance, utilizing a spectrum from blue (lower scores) to red (higher scores). Edge thickness corresponds to interaction frequency between student pairs.
(a) Explain Own Idea Network, Pre-test
(b) Explain Own Idea Network, Post-test
(c) Engage Others’ Idea Network, Pre-test
(d) Engage Others’ Idea Network, Post-test
Figure 2: Explain Own Idea and Engage Others’ Idea Networks. Edge weights in each network represent the frequency of corresponding utterances, and node colors represent math test scores (red = higher, blue = lower). The left-hand figures show pre-test results, and the right-hand figures show post-test results. The networks are constructed using data aggregated from six lessons.
The network visualization in Figure 2 reveals that interaction quality, rather than quantity, plays a more significant role in influencing learning outcomes. For instance, Julian and William, who initially demonstrated lower mathematics performance in the pre-test, established strong connections primarily with high-achieving peers within the network. Their subsequent improvement in post-test scores suggests that interactions with high-performing students may create multiplicative learning benefits. Conversely, Mya’s case provides an instructive contrast: despite maintaining frequent peer interactions, her academic performance showed minimal improvement. This pattern reinforces the observation that the qualitative aspects of student interactions — specifically, the academic performance levels of interaction partners — may exert greater influence on learning outcomes than the mere frequency of exchanges.
Network Centrality
We then examined network centrality, using metrics designed for directed weighted networks, accounting for both directional flow and interaction intensity. These include PageRank (Page et al., 1999), hub-authority scores (Kleinberg, 1999), in-degree and out-degree, betweenness (Freeman, 1977), closeness (Bavelas, 1948), and eigenvalue centrality (Bonacich, 1972).
Each centrality measure captures distinct aspects of student engagement in the network. PageRank evaluates student influence by considering both the number and importance of incoming connections, assigning higher scores to students who receive interactions from influential peers. Hub scores identify students who initiate connections with multiple peers, while authority scores highlight students who serve as primary information sources. In-degree and out-degree centrality measure the weighted volume of received and initiated interactions respectively, identifying key participants in collaborative discourse. Eigenvalue centrality assesses each student’s influence by considering the influence levels of their connected peers.
Betweenness and closeness centrality measures require a different approach to weight calculation, as they interpret edge weights as distances rather than strengths. We employed the inverse of utterance frequencies as weights, ensuring that frequent interactions correspond to shorter distances. Consequently, betweenness centrality identifies students who serve as key intermediaries in information flow, while closeness centrality quantifies each student’s overall proximity to peers in the interaction network. This inverse weighting scheme ensures that stronger connections (more frequent interactions) contribute to higher centrality scores.
  Pagerank Indegree Outdegree Betweenness Closeness Eigen Hub Authority
  Angelina 0.1033 57 61 32 0.3695 0.2245 0.0378 0.0654
Armando 0.0414 46 46 34 0.4829 0.2242 0.1974 0.0198
Daniel 0.0286 48 37 27 0.4728 0.1811 0.0216 0.6029
Diego 0.0451 75 47 115 0.5616 0.5980 0.0147 0.7355
Esther 0.0479 92 105 76 0.5176 0.8422 0.0556 0.6152
Gio 0.0515 56 139 145 0.5454 0.2639 1.0000 0.0171
Isabel 0.0330 42 40 29 0.4775 0.2533 0.2659 0.0570
Jessica 0.0522 71 26 57 0.4916 0.7276 0.0133 0.5414
Julian 0.0530 87 51 52 0.5046 0.3090 0.0385 0.9111
Keila 0.0229 35 22 0 0.3728 0.1329 0.0080 0.0877
Kevin 0.0484 89 55 71 0.4571 0.2615 0.0045 1.0000
Kimberly 0.0481 51 93 51 0.4613 0.2644 0.1562 0.0420
Marco 0.0518 49 73 18 0.3276 0.1015 0.0067 0.2131
Maria 0.0607 41 53 61 0.4259 0.2476 0.1405 0.0570
Melissa 0.0184 25 46 0 0.3872 0.2847 0.1741 0.0217
Mya 0.0665 92 135 142 0.5677 1.0000 0.6806 0.0618
Natalie 0.0343 60 37 4 0.4005 0.4354 0.0421 0.2712
Ruben 0.0481 59 74 43 0.4086 0.1663 0.5060 0.0051
Samantha 0.0965 77 38 5 0.3629 0.2925 0.0247 0.0741
William 0.0485 62 36 7 0.3242 0.1290 0.1091 0.0133
Table 6: Centrality measures for the EXP (Explain Own Idea) network.
  Pagerank Indegree Outdegree Betweenness Closeness Eigen Hub Authority
  Angelina 0.0307 48 48 36 0.2968 0.0045 0.0006 0.0030
Armando 0.0431 96 50 9 0.6094 0.4061 0.0796 0.0262
Daniel 0.0358 36 33 0 0.4462 0.1607 0.0215 0.0174
Diego 0.0556 96 128 135 0.6766 0.5410 0.0292 0.4539
Esther 0.0964 192 75 110 0.6310 1.0000 0.0162 1.0000
Gio 0.0744 134 75 125 0.5837 0.3982 0.0477 0.0443
Isabel 0.0377 73 38 17 0.5011 0.2453 0.0999 0.0430
Jessica 0.0278 59 80 90 0.5569 0.1406 0.0098 0.1212
Julian 0.0600 71 152 57 0.5309 0.2287 0.0521 0.0290
Keila 0.0228 33 40 0 0.4228 0.1101 0.0055 0.0623
Kevin 0.0615 75 86 38 0.4894 0.2484 0.0398 0.0203
Kimberly 0.0537 73 97 63 0.5287 0.2807 0.2482 0.0173
Marco 0.0707 104 54 10 0.4116 0.2906 0.0092 0.0144
Maria 0.0275 51 77 68 0.5223 0.0788 0.0449 0.0060
Melissa 0.0481 59 58 11 0.5133 0.3470 0.3616 0.0274
Mya 0.0646 119 175 167 0.6586 0.6136 1.0000 0.0229
Natalie 0.0329 47 67 0 0.4311 0.1986 0.0909 0.1434
Ruben 0.0629 82 76 29 0.4609 0.2486 0.0209 0.0447
Samantha 0.0276 37 47 0 0.2647 0.0019 0.0044 0.0002
William 0.0662 71 100 20 0.4313 0.2460 0.0153 0.0144
Table 7: Centrality measures for the EOI (Engage Others’ Idea) network.
Tables 6 and 7 illustrate how different centrality measures reveal distinct patterns of student engagement in the interaction networks. For example, in the explanation network, Mya demonstrates high PageRank centrality, indicating her frequent interactions with influential peers. Esther’s elevated in-degree and out-degree scores reflect her active participation as both recipient and initiator of mathematical explanations. Notably, Gio’s high betweenness centrality suggests a crucial role in facilitating information flow between different student groups, effectively bridging distinct classroom communities.
Analysis of Table 7 reveals distinct interaction patterns within the Engagement network, with several students demonstrating notable centrality profiles. Esther’s consistently high scores across centrality metrics indicate her comprehensive engagement in mathematical discourse, both influencing and responding to peers’ ideas. Gio maintains his bridging role while actively participating in peer discussions, while Mya demonstrates strong bilateral engagement through both contributing to and building upon others’ mathematical reasoning. Additionally, Armando, Diego, and Julian’s elevated eigenvector centrality and hub scores suggest their significant influence within localized peer groups, indicating the presence of distinct collaborative clusters within the classroom.
Our analysis shows that high network centrality does not necessarily correlate with improved mathematical performance. Mya’s case illustrates this complexity: despite her prominent position in both interaction networks, her mathematical achievement did not improve, possibly due to her frequent interactions with lower-performing peers. This suggests that the relationship between network centrality and academic performance is not straightforward and the specific composition of local interactions - such as which students interact with whom - appears to play a more significant role. The influence of classroom interactions on learning outcomes appears to depend more on the composition of immediate peer groups than on students’ overall network prominence. This observation underscores the importance of examining both the quality and context of student interactions within local network structures, rather than relying exclusively on global centrality metrics.
3.2 Network Mediation Model
Following the descriptive analysis of classroom interaction networks through visualization and centrality metrics, we now proceed to a more rigorous examination of how these network structures influence student outcomes. While network centrality measures offer valuable insights into the roles and prominence of individual students within the interaction network, they do not directly reveal the mechanisms through which these interaction patterns impact learning. To understand these mechanisms, we employ network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022). Network mediation analysis provides a framework for addressing this challenge by treating the network structure as a mediating variable. A key challenge in implementing network mediation analysis is representing the complex network structure in a way that can be used as a mediator. To address this, we adopt the approach of latent space model (Hoff et al., 2002; Hoff, 2005; Sewell and Chen, 2016, LSM)). LSM provide a framework to represent each student’s position within the network in a lower-dimensional latent space, capturing their underlying social dynamics wherein inter-student distances reflect the probability and intensity of their interactions.
Figure 3: Multiple mediator model for network mediation analysis
Figure 3 presents a diagram of a multiple mediator model for network mediation analysis. In this model, $M_{1}$ through $M_{k}$ represent the latent positions of network nodes embedded in a latent space. The paths $a_{11}$ through $a_{1k}$ denote the effects of $X$ on the mediators, while the paths $b_{21}$ through $b_{2k}$ capture the effects of the mediators on $Y$. The direct effect of $X$ on $Y$ is represented by path $b_{1}$. By linking these latent positions to independent and dependent variables through regression frameworks, we can decompose total effects into direct and indirect components mediated by the underlying network structure.
Network as a Mediator for Gender Gap in Mathematics Performance
Our study examines how classroom interaction networks mediate the influence of gender on mathematics achievement. Prior research has consistently investigated gender differences in mathematics performance, revealing complex patterns and influences (Lindberg et al., 2010; Else-Quest et al., 2010; Ghasemi and Burley, 2019), while Lindberg et al. (2010) demonstrate that overall gender gaps in mathematics performance are relatively small. Ghasemi and Burley (2019) identifies significant cultural and societal influences on mathematics-related preferences and self-confidence across genders. Else-Quest et al. (2010) emphasizes the role of societal context in gender disparities, noting smaller performance differences in countries with greater gender equality. These findings collectively suggest that gender differences in mathematics achievement stem from social rather than purely biological factors, underscoring the importance of analyzing classroom social networks and their mediating effects on academic performance.
Based on this perspective, we hypothesize that the gender gap in mathematics performance ($X$ on $Y$) can be differentiated based on how individual students interact with others during mathematics discussion classes. Specifically, we apply network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022), using students’ interaction structures as a mediator ($M$) for gender ($X$) - math performance ($Y$) relationships. By incorporating the estimated latent positions ($z_{i}$, $w_{j}$) as mediators, our goal is to investigate the indirect effects of gender on mathematics performance through the network structure.
Pre-test Scores as a Confounder
In the current context, pre-test scores could serve as a useful confounder for the intended mediation analysis. However, the pre-test (California Standardized test; CST) and post-test (a test developed by Webb et al. (2023)) do not measure the same domains, while the assessment methods and criteria are also different between the two tests. Thus, the two test scores may not be entirely comparable (Liu et al., 2017).
Figure 4: Scatter plot showing the relationship between pre-test (CST) and post-test scores.
Further, there is an additional technical challenge: the pre-test scores do not seem to have a linear relationship with the post-test scores. As shown in Figure 4, the relationship is not clearly linear around the Basic (300-350) and Proficient (350-400) ranges of the pre-test. This means that higher scores in the pre-test may not predict higher post-test scores consistently, suggesting pre-test scores may not be a reliable predictor of the post-test scores.
To account for students’ prior mathematical proficiency in our analysis, we constructed a binary proxy variable based on California Standards Test (CST) performance. Students who achieved scores of 350 or above — the state-designated threshold for proficiency [^2] https://www.starsamplequestions.org/about.html — were categorized as high-performing. This dichotomization aligns with state educational standards and provides a meaningful distinction in student achievement levels. We incorporate this proxy variable as a confounder in our mediation analysis to more precisely evaluate how network structures mediate gender-based differences in mathematics performance.
Figure 5: Network mediation model path diagram. This model analyzes the effect of gender on math achievement mediated by network interactions. Latent positions of sender and receiver networks are used as mediators.
Network Mediation Model
We additionally consider predictor-mediator interaction effects to account for potential heterogeneity in mediator effects across varying predictor levels (VanderWeele and Vansteelandt, 2009; Valeri and VanderWeele, 2013). In this case, the indirect effects, mediated through the latent network structure, vary with gender ($x$). By modeling this interaction, we account for the possibility that the indirect effect of gender on math performance is mediated differently depending on the students’ positions within the network. The resulting network mediation model can be expressed as:
$\begin{split}E(M_{k}\mid x,c)&=a_{0k}+a_{1k}\,x+a_{2k}\,c,\\ E(Y\mid x,\mathbf{m},c)&=b_{0}+b_{1}\,x+\sum_{k}b_{2k}\,m_{k}+\sum_{k}b_{3k}\,x\,m_{k}+b_{4}\,c,\end{split}$ (1)
where $x$ is the binary indicator for gender, $m_{k}$ represents the latent positions of students, and $c$ is the confounder indicating proficiency based on the CST score threshold. Specifically, the latent variables $m_{k}$ represent the coordinates of the latent positions for the network nodes: $m_{1}$ and $m_{2}$ correspond to sender position coordinates (x and y axes, respectively), while $m_{3}$ and $m_{4}$ represent receiver position coordinates (x and y axes, respectively). The model coefficients capture distinct aspects of the gender-performance relationship: $a_{1k}$ quantifies gender’s influence on network positioning, $b_{2k}$ measures how network positions affect mathematics performance, and $b_{3k}$ captures the predictor-mediator interaction between gender and network position, revealing how network-mediated effects vary by gender. Additionally, $a_{2k}$ and $b_{4}$ account for the influence of prior mathematical proficiency on network positions and post-test performance, respectively. $a_{0k}$ and $b_{0}$ are intercept terms for the network mediation model.
This model enables the decomposition of the total effect (TE) into two distinct components: the natural direct effect (NDE) and the natural indirect effect (NIE). The NDE represents the component of the effect of gender ($X$) on math performance ($Y$) that is independent of the mediators, while the NIE captures the effect transmitted through the mediating network structure. These components constitute the total effect (TE), providing a comprehensive measure of the impact of gender on math performance. The natural direct effect, the natural indirect effect, and the total effect, are given by:
$\begin{split}NDE&=\left(b_{1}+\sum_{k}b_{3k}\left(a_{0k}+a_{1k}\,x^{*}+a_{2k}\,c\right)\right)(x-x^{*}),\\ NIE&=\left(\sum_{k}\left(b_{2k}\,a_{1k}+b_{3k}\,a_{1k}\,x\right)\right)(x-x^{*}),\\ TE&=NDE+NIE.\end{split}$ (2)
These effects can be calculated from the regression coefficients estimated with Equation (1).
Our analysis examines how distinct types of classroom interactions mediate gender-based differences in mathematics performance by investigating two separate networks: the EXP network and the EOI network. For both networks, we incorporate students’ prior mathematics achievement as a confounder to account for pre-existing academic differences. Through the application of our network mediation model to these distinct interaction networks, we aim to investigate whether the mediating effects of classroom network structures vary systematically based on interaction type.
3.3 AMEN Model for Latent Position Estimation
To obtain the latent positions that serve as mediators, we employ the AMEN model (Hoff, 2021). The AMEN model is an extension of latent space model (LSM; Hoff et al., 2002), designed to capture more sophisticated network patterns, particularly in directed networks like classroom interactions. This extension is particularly relevant for our analysis of classroom interactions, which are inherently directed. The AMEN model explicitly models directed relationships by providing distinct latent positions for each student as both a sender and a receiver. The AMEN model can be expressed as:
$g(\mathbb{E}(y_{ij}))=\alpha_{i}+\beta_{j}+z_{i}^{T}w_{j},$ (3)
where $y_{ij}$ is the value representing the connection between the $i$th node and the $j$th node (e.g., the number of interactions, a binary indicator of whether an interaction occurred), $\alpha_{i}$ and $\beta_{j}$ represent sender and receiver random effects, respectively, and $z_{i}$ and $w_{j}$ are low-dimensional latent feature vectors for nodes $i$ as a sender and $j$ as a receiver respectively, in a $d$-dimensional latent space, where $d=2,3$ is typically considered for visualization purposes. $g(\cdot)$ is a link function. The additive components, $\alpha_{i}$ and $\beta_{j}$, account for individual node-specific effects, while the multiplicative term, $z_{i}^{T}w_{j}$, captures interactions between latent vectors. This approach allows the model to represent complex third-order network dependencies, such as transitivity, clustering, and degree heterogeneity.
AMEN with a Negative Binomial Regression
A unique feature of classroom interaction data is its count-based nature, reflecting the frequency of student interactions. By incorporating these interaction frequencies as edge weights in the network structure, we intend to capture both the presence and intensity of student-to-student connections.
For count data, a Poisson link function is typically considered. However, a Poisson model often suffers from overdispersion, where the observed variance exceeds the theoretical mean. The sender and receiver random effects can account for overdispersion to some degree by allowing the conditional mean to vary across participants (Gelman, 2007). However, severe overdispersion may not be fully addressable with the sender and receiver random effects.
In our analysis of classroom interaction data, we identified a severe overdispersion issue, as detailed in Section D of the Supplementary Material. To effectively address overdispersion, we propose using a negative binomial regression model, which introduces a dispersion parameter to explicitly account for the excess variability and allows the conditional variance to vary independently of the mean. By combining the negative binomial specification with sender and receiver random effects from the AMEM model, we can effectively capture both individual-level heterogeneity and the overdispersed nature of classroom interaction counts.
The proposed AMEN model with a negative binomial regression structure for count data can be written as:
$\log(\mu_{ij})=\alpha_{i}+\beta_{j}+z_{i}^{T}w_{j},$ (4)
where $\mu_{ij}$ is the mean of the negative binomial distribution where $y_{ij}\sim\mbox{NB}(\mu_{ij},r)$ and $r$ of the negative binomial distribution controls the variance of the distribution, with higher values indicating greater dispersion relative to the mean. $\alpha_{i}$ and $\beta_{j}$ are sender and receiver random effects, respectively, and $z_{i}$ and $w_{j}$ are sender and receiver latent feature vectors, respectively.
3.4 Estimation Methods
We take a two-stage estimation approach for data analysis: (1) AMEN model estimation; and (2) network mediation model estimation.
AMEN Model Estimation
In the first stage, we estimate the proposed AMEN model to embed the observed student interaction network into a latent space. The prior distributions of the AMEN model parameters are specified as follows:
$\begin{split}\alpha_{i}&\sim N(0,\sigma_{\alpha}^{2}),\quad\sigma_{\alpha}^{2}>0,\quad i=1,\dots,N\\ \beta_{j}&\sim N(0,\sigma_{\beta}^{2}),\quad\sigma_{\beta}^{2}>0,\quad j=1,\dots,N\\ z_{i}&\sim N_{d}(0,\sigma_{z}^{2}I_{d}),\quad\sigma_{z}^{2}>0,\quad i=1,\dots,N\\ w_{j}&\sim N_{d}(0,\sigma_{w}^{2}I_{d}),\quad\sigma_{w}^{2}>0,\quad j=1,\dots,N\\ \log r&\sim N(\mu_{r},\sigma_{r}^{2}),\quad\mu_{r}\in\mathbb{R},\quad\sigma_{r}^{2}>0\end{split}$ (5)
where $\sigma_{\alpha}^{2}$, $\sigma_{\beta}^{2}$, $\sigma_{z}^{2}$, $\sigma_{w}^{2}$, and $\sigma_{r}^{2}$ represent fixed prior variance terms for each corresponding parameter, and $I_{d}$ denotes the $d\times d$ identity matrix. In this study, we set $d=2$ for interpretability and visualization purposes.
We draw the posterior samples of the AMEN model parameters and latent positions using a Metropolis-Hastings-within-Gibbs sampler. Additional details on the prior means, standard deviations, and the jumping rule for each parameter, as well as the starting points and overall convergence checks, are provided in the Section A of the Supplementary Material.
Note that the inner product between latent vectors remains invariant under transformations such as rotation or reflection. To address this, we perform Procrustes matching (Gower, 1975) to align the latent vectors to a common reference position. This alignment ensures that the latent vectors are consistently positioned, facilitating a clearer interpretation of their relationships and enhancing the validity of subsequent analyses.
Network Mediation Model Estimation
In the second stage, we estimate a network mediation model. Specifically, the latent positions from the estimated AMEN model serve as mediators in the regression framework outlined in Equation (2). The network mediation model is estimated through Bayesian regression, with prior distributions specified as:
[…]
where $\Theta\in\{a_{01},\dots,a_{2k},b_{0},b_{1},b_{21}\dots,b_{3k},b_{4}\}$ represents all regression coefficients in Equation (1), $\epsilon_{i}$ represents the error term in the regression model, and $\sigma^{2}$ is the variance of error term. The network mediation model is estimated using the MCMCpack package (Martin et al., 2011) in R. The estimated regression coefficients are then used to calculate the natural direct effect (NDE) and natural indirect effect (NIE), as defined in Equation (2).
3.5 Network Analysis Results
3.5.1 AMEN Model Results
(a) EXP (Explain Own Idea) Network
(b) EOI (Engage Others’ Idea) Network
Figure 6: Comparison of Latent Positions for EXP and EOI Networks. Each node is represented as a point in a two-dimensional space, where the direction from the origin indicates similarity between nodes, and the distance from the origin indicates influence within the network.
Latent Position Structure
Figures 6(a) and 6(b) illustrate the latent positions of students within both the EXP and EOI networks through two-dimensional visualizations. In these representations, each student’s position is mapped as a point in two-dimensional space. The angle between student vectors indicates their interaction similarity - smaller angles represent greater similarity in interaction patterns. Additionally, vector magnitude represents each student’s relative influence within the network structure.
Figure 2(a) reveals important bridging roles within the network structure, particularly through the positions of Gio and Isabel. Gio functions as a critical connector between two distinct student subgroups: one comprising Diego, Daniel, and Armando, and another consisting of Julian, William, and Marco. Similarly, Isabel facilitates connections between the subgroup of Melissa, Natalie, and Esther and the subgroup of Julian, William, and Marco. The latent vector visualization in Figure 6(a) captures these bridging relationships effectively, positioning Gio centrally between the connected subgroups while placing Isabel in closer proximity to the subgroup of Melissa, Natalie, and Esther.
In Figure 2(c), Isabel serves a bridging role in connecting two distinct student clusters: the Keila and Kimberly group and the Marco and Julian group. In Figure 6(b), Isabel appears closer to the Keila, Kimberly, Melissa, Natalie, and Esther groups, reflecting the higher frequency of interactions within this group compared to the Marco and Julian cluster. The visualization also captures subtle relationship dynamics, such as Esther’s unidirectional engagement with William, which influences the latent space positioning of both Isabel and William through second-order interaction effects. These examples demonstrate the advantages of latent vector analysis in visualizing complex network data. By projecting interactions into a low-dimensional space, this approach uncovers hidden structures, including higher-order interactions and bridging relationships, while precisely representing both directionality and intensity of student connections. This methodology provides valuable insights into group dynamics, particularly for applications in educational research.
3.5.2 Network Mediation Model Results
Our investigation extends to examining how interaction patterns influence students’ mathematical achievement through network mediation analysis. This analytical approach specifically evaluates how network structures mediate the relationship between gender and mathematics learning outcomes. The network mediation analysis quantifies the three key components: the natural indirect effect (NIE), the natural direct effect (NDE), and the total effect (TE). Tables 8 and 9 present the comprehensive results of these analyses.
Post.M Post.SD 2.5% 97.5%
NIE 0.7758 1.2468 -1.7182 3.3649
NDE -5.475 1.5948 -8.8699 -2.48 *
TE -4.6992 0.8321 -6.4324 -3.056 *
Table 8: Network Mediation Model Results for EOI Network
Analysis of the EOI network, presented in Table 8, indicates that NDE is statistically significant and negative, while NIE through the network is not statistically significant. This negative NDE after controlling for proficiency of pre-test as a confounder, suggests that gender has a clear negative direct effect on post-test scores, meaning female students tend to achieve lower math scores independent of the mediating EOI network structure.
Post.M Post.SD 2.5% 97.5%
NIE 2.8653 1.3964 0.5089 5.9881 *
NDE -6.9407 1.9286 -11.0408 -3.3803 *
TE -4.0754 1.1155 -6.344 -1.7559 *
Table 9: Network Mediation Model Results for EXP Network
In contrast, for the EXP network, Table 9 reveals a different pattern. The negative NDE, which is consistent with the EOI network, shows that female students’ scores tend to be lower when considering gender independently of network effects. However, the positive NIE suggests that gender influences network structure in ways that enhance female students’ performance. These findings highlight the crucial role of network interactions, particularly within the EXP network, in mediating and potentially mitigating gender-based performance differences through collaborative learning processes.
4 Discussion
In the current study, we proposed the analysis of classroom interactions from their dialogues through an innovative combination of prompt engineering and network analysis methodologies. The prompt engineering approach enables automated classification of interaction patterns without requiring extensive pre-labeled datasets, significantly improving upon traditional manual labeling methods in both efficiency and scalability. Building on this automated classification, our network analysis framework represents students as nodes and their interactions as edges, providing a comprehensive visualization and analysis of classroom information flow and engagement patterns. Through the projection of these networks onto latent spaces, we quantified structural characteristics that provide the understanding structure of classroom interactions.
The network mediation analysis provides important insights into the relationship between classroom dynamics and student achievement. Our findings demonstrate that network mediation models possess unique capabilities for capturing intricate social patterns in educational environments which traditional analytical approaches often fail to identify due to the multidimensional nature of classroom interactions. This research opens several promising directions for future research, particularly in the incorporation of additional variables and the development of more sophisticated network representations, such as dynamic networks to capture temporal evolution of interactions or implement multilayer networks to represent diverse forms of student engagement simultaneously. Our methodological integration validates the effectiveness of combining prompt engineering with network analysis for understanding how student interactions influence learning outcomes.
Our findings suggest several promising directions for future research to deepen our understanding of classroom dynamics and refine analytical methodologies. A key area for expansion involves the analysis of teacher-student interactions, which we did not address in the current study. Teachers serve as essential facilitators of learning, significantly influencing classroom dynamics and student engagement patterns. Examining these teacher-student interactions would provide valuable insights into how instructional practices shape educational outcomes, particularly in the development of critical thinking skills and collaborative learning. Such analysis could inform the optimization of teaching strategies and contribute to enhanced classroom effectiveness.
Future research could enhance the accuracy of interaction analysis by exploring model fine-tuning with high-quality labeled datasets, extending beyond our current prompt engineering approach. Such fine-tuning could significantly improve the model’s precision and generalizability in analyzing complex educational interactions. Additionally, the systematic integration of expert knowledge and carefully curated datasets into the prompt engineering process would strengthen the methodological reliability and validity. This enhanced approach could provide more nuanced insights into classroom interaction patterns while maintaining the efficiency advantages of automated analysis.
A critical direction for future research involves developing automated prompt-generation methodologies. The current reliance on human judgment for prompt creation and selection introduces potential variability and bias into the analytical process. An automated approach that selects prompts based on their likelihood of generating desired outputs would enhance both the efficiency and consistency of the labeling process. Such automation would transform this methodology into a more scalable and reliable tool for educational research, enabling broader application across diverse educational contexts while maintaining analytical rigor.
Acknowledgement
We thank the editor, associate editor, and reviewers for their constructive comments. This work was partially supported by the National Research Foundation of Korea [grant number RS-2023-00217705 and RS-2024-00333701; Basic Science Research Program awarded to IHJ] and the ICAN (ICT Challenge and Advanced Network of HRD) support program [grant number RS-2023-00259934], supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation). Correspondence should be addressed to Ick Hoon Jin, Department of Applied Statistics, Department of Statistics and Data Science, Yonsei University, Seoul, Republic of Korea. E-mail: ijin@yonsei.ac.kr.

### v2

@@
 1.1 Background: Classroom Dialogue Analysis
-Classroom interactions play a crucial role in the learning process and substantially influence students’ motivation, comprehension, and academic achievement.
-Through effective student-teacher and student-student interactions, learners develop essential skills in critical thinking, collaborative work, and problem-solving.
-Analysis of these student interactions reveals underlying behavioral patterns that either facilitate or impede effective learning.
+Classroom interactions play a crucial role in the learning process and substantially influence student motivation, comprehension, and academic achievement.
+Through effective student-teacher and student-student interactions, learners develop essential skills in critical thinking, collaborative work, and problem solving.
+Analysis of these student interactions reveals underlying behavioral patterns that facilitate or impede effective learning.
 The analysis of classroom dialogues provides a powerful lens for examining student interactions, offering a direct understanding of engagement patterns and learning processes.
@@
 These conversations are often characterized by dynamic shifts in topic, speaker turns, and the interplay of various communicative functions, making manual analysis particularly difficult.
-These also encompass speaker characteristics (e.g., their speaking style, experience), emotional tone (e.g., enthusiasm, uncertainty, affect), and context (e.g., type of task, classroom environment).
+They also encompass speaker characteristics (e.g., their speaking style, experience), emotional tone (e.g., enthusiasm, uncertainty, affect), and context (e.g., type of task, classroom environment).
 To address such complexity, researchers often simplify the analysis by categorizing the intent of individual messages or utterances using predetermined coding schemes.
@@
 […]
-Strategic prompt engineering enables researchers to classify dialogue interaction types even without traditional training data for fine-tuning, provided the prompts incorporate sufficient contextual information.
-Nevertheless, while prompt engineering facilitates classification of individual utterances, it cannot fully capture the complex dynamics of classroom interactions.
-The classification of discrete utterances reveals speaker identities and discussion content but fails to illuminate critical patterns such as interaction frequency, structural relationships, and the relationships between students.
+Building on these capabilities, recent studies have explored prompt engineering for automating classroom dialogue analysis.
+Long et al.
+(2024) reported that LLMs like GPT-4 can perform coding tasks up to 30 times faster than human researchers while achieving over 90% agreement for relatively straightforward classification tasks.
+On the other hand, Ganesh et al.
+(2024) found that while models perform well on tasks with clear linguistic cues, they exhibit unstable and inconsistent performance on complex, theoretically motivated qualitative coding tasks requiring nuanced interpretation.
+Ganesh et al.
+(2024) added that fine-tuned deep learning models (i.e., models with parameters updated using labeled training data) often outperform prompt-based approaches, which rely solely on prompt design without parameter updates.
+These findings highlight the need for methodological innovations that can enhance prompt-based classification reliability while preserving the flexibility advantages of not requiring extensive pre-labeled datasets.
+Despite these challenges, strategic prompt engineering still enables researchers to classify dialogue interaction types without traditional training data for fine-tuning, provided the prompts incorporate sufficient contextual information.
+Nevertheless, even with enhanced prompt engineering approaches, a fundamental limitation remains: while prompt engineering facilitates classification of individual utterances, it cannot fully capture the complex relational dynamics of classroom interactions.
+The classification of discrete utterances reveals speaker identities and discussion content but fails to illuminate critical patterns such as interaction frequency, structural relationships, and relationships among students.
 Network Analysis
@@
 In this paper, we present an integrated framework for analyzing classroom interactions by combining the strengths of prompt engineering and network analysis.
-Our methodology begins with automated dialogue classification through prompt engineering, addressing the limitations of traditional manual labeling approaches.
+Our methodology begins with automated utterance classification through prompt engineering, addressing the limitations of traditional manual labeling approaches.
 We employ carefully constructed prompts to guide the generative model in classifying interactions according to predefined criteria, establishing the foundation for network-based analysis of classroom dynamics.
+To enhance classification reliability, we employ an ensemble strategy that aggregates the predictions of multiple instruction-tuned LLMs using majority voting.
+This simple yet effective method produces more consistent labels (Yue et al., 2024) and helps mitigate limitations of individual models, such as biases, sensitivity to prompts, and task-specific variability (Jiang et al., 2023; Lu et al., 2024; Chen et al., 2025).
+By building consensus across models, the ensemble approach enables more robust utterance classification and improves the reliability of subsequent network analyses compared to relying on a single model.
+Further, we incorporate an uncertainty quantification framework to systematically detect ambiguous utterances that may benefit from human review.
+This involves assessing disagreement among models and evaluating token-level probability distributions to gauge labeling uncertainty.
+Utterances with high entropy or inconsistent probability patterns across models are flagged for expert inspection.
+In this way, annotation efforts can be allocated more efficiently while ensuring strong reliability and accuracy in classification.
 We then transform these classified interactions into network representations and analyze them using a network mediation model.
-Specifically, we employ the additive and multiplicative effects network (AMEN) model (Hoff, 2021) model to extract latent social structures and utilize them as mediators to investigate how classroom networks mediate the relationship between gender and mathematics performance.
+Specifically, we employ the additive and multiplicative effects network model (AMEN) (Hoff, 2021) to extract latent social structures and use them as mediators to investigate how classroom networks mediate the relationship between gender and mathematics performance.
 Through examination of how these latent social structures influence student outcomes, with particular attention to gender differences in mathematics performance, we illuminate the mechanisms by which network structures mediate the impact of classroom dynamics on student achievement.
-This comprehensive methodological integration can provide educational researchers with a powerful analytical tool for investigating the complex relationships between social dynamics and academic performance.
+This comprehensive methodological integration can provide educational researchers with a powerful analytical tool to investigate the complex relationships between social dynamics and academic performance.
 The remainder of this paper is structured to present our methodology and findings systematically.
@@
 Student A: Yeah.
-Student C: Mine got 35.
+Student C: Mine got 35. I think you missed one group.
 […]
 2.2 Prompt Engineering Methodology
-The complexity and high-dimensional nature of classroom interactions necessitate a robust method for classifying and analyzing dialogue intent.
-As discussed in the Introduction Section, existing analytical approaches require manual labeling, which is resource-intensive, while deep learning methods require extensive labeled datasets.
+The complexity and high-dimensional nature of classroom interactions require a robust method to classify and analyze the intent of dialogue.
+As discussed in Section 1, existing analytical approaches require manual labeling, which is resource-intensive, while deep learning methods require extensive labeled datasets.
 Our methodology addresses these issues through prompt engineering, by leveraging pre-trained models to classify dialogue intents accurately without labeled training data or computational fine-tuning.
 This approach maximizes the capabilities of pre-trained models while maintaining efficiency in resource utilization and processing time.
-In-context learning (ICL) is a specialized application of prompt engineering in which pre-trained large language models (LLMs) learn exclusively from information provided within the prompt, eliminating the need for parameter updates.
-This approach parallels human learning processes through inference while minimizing computational requirements for task adaptation (Dong et al., 2022).
+In-context learning (ICL) is a specialized application of prompt engineering in which pre-trained LLMs learn exclusively from information provided within the prompt, eliminating the need for parameter updates.
+This approach parallels human learning processes through inference while minimizing computational requirements for task adaptation (Dong et al., 2024).
 Few-shot prompting, one of the ICL strategies, enhances model comprehension by incorporating targeted task examples within the prompt.
-The effectiveness of few-shot prompts in ICL has been explained through probabilistic frameworks such as hidden Markov models (Baum and Petrie, 1966), where carefully selected examples guide model reasoning by establishing sequential dependencies that reduce prediction uncertainty (Xie et al., 2021).
+The effectiveness of a few-shot prompts in ICL has been explained through probabilistic frameworks such as hidden Markov models (Baum and Petrie, 1966), where carefully selected examples guide model reasoning by establishing sequential dependencies that reduce prediction uncertainty (Xie et al., 2021).
 Min et al.
-(2022) further demonstrates that optimal performance depends not merely on increasing example quantity but on developing prompt formats that precisely align with task-specific requirements, including both structural elements and linguistic formulation.
+(2022) further demonstrates that optimal performance depends not merely on increasing the amount of examples, but on developing prompt formats that precisely align with the specific requirements of the task, including both structural elements and linguistic formulation.
+In our study, the prompts are systematically designed using a two-step chain-of-thought structure to improve classification accuracy and consistency.
+The first step generates explicit reasoning for the classification decision, while the second step applies this reasoning to produce the actual categorization.
+The system prompt incorporates comprehensive classification criteria and categorization rules that define the utterance classification task.
+Specifically, it provides explicit guidelines for categorizing interactions as “explaining own ideas” or “engaging with others’ ideas,” with the latter further differentiated into low, medium, or high engagement levels.
+In addition, the system prompt outlines specific criteria for making classification decisions, including handling special cases and exceptions.
+Throughout prompt design, we emphasize the importance of accurate categorization and consistency across classifications to ensure that subtle distinctions between labels are properly recognized and maintained.
+The user prompt complements this framework by providing essential contextual information required for classification decisions, including previous utterances and speaker information that enable the model to understand the flow of conversation and the relationships between participants.
+This contextual grounding facilitates more nuanced classification of utterances into the three main categories: explain your own ideas, engage with others’ ideas, or uncorrelated.
+The prompts incorporate illustrative dialogues that clarify distinctions between categories and demonstrate practical application of classification rules.
+Detailed prompts are provided in Section LABEL:supp-sec:prompt of the Supplementary Material.
+To enhance output reliability and reduce variability in model responses, we configure the temperature parameter to zero, ensuring reproducible results for our analysis.
+To enhance classification robustness and reduce individual model biases, we employ an ensemble of seven instruction-tuned LLMs comprising both open-source models (Meta’s Llama-3.1-8B (Dubey et al., 2024), Alibaba’s Qwen-3-32B-AWQ (Yang et al., 2025), and Microsoft’s Phi-4 (Abdin et al., 2024)) and commercial models (OpenAI’s GPT-4.1 (OpenAI, 2025), Google’s Gemini-2.5-Pro (Google DeepMind, 2025a), Gemini-2.5-Flash (Google DeepMind, 2025b), and Anthropic’s Claude-4-Sonnet (Anthropic, 2025)).
+These models exhibit distinct response patterns due to variations in training data, fine-tuning approaches such as reinforcement learning from human feedback and supervised tuning, and underlying development methodologies.
+Rather than relying on predictions from a single model, we collect responses from all ensemble members and assign the most frequently occurring classification as the final label through majority voting.
+This ensemble approach mitigates individual model biases and produces more robust classification labels through cross-model consensus.
+All models are prompted using a standardized few-shot format that incorporates task-specific interaction examples designed to improve instruction adherence and classification accuracy.
+Prompt construction adheres to the official formatting guidelines for each model to ensure methodological consistency and reproducibility across the entire ensemble.
+Detailed model descriptions and formatting guidelines are provided in Section LABEL:supp-sec:llm of the Supplementary Material.
+This standardized prompting approach ensures equitable comparison and consistent behavioral patterns across all ensemble members, effectively mitigating the influence of model-specific idiosyncrasies on classification outcomes.
+2.3 Utterance Classification Results
+The classification process implements a two-step chain-of-thought structure, requiring models to first generate explicit reasoning before producing the final categorization.
+Each few-shot example demonstrates comprehensive reasoning processes covering speaker identification, utterance content analysis, classification labeling, engagement level assessment, and relevant contextual references.
+Complete prompt specifications, including few-shot examples, are provided in Section LABEL:supp-sec:prompt of the Supplementary Material.
+The classification process was executed on an NVIDIA RTX A6000 GPU with an Intel Xeon Gold 6426Y CPU in a Linux environment.
+Following the utterance classification, we excluded entries labeled as “uncorrelated” from subsequent analyses.
+The remaining utterances, classified as EXP and EOI, were segregated to construct distinct interaction networks for each type of discourse.
+As a result, a total of 4,233 utterances were analyzed, 497 classified as EXP and 696 as EOI.
+The latter category was further divided into three subcategories based on engagement level: high, medium, and low.
+Table 1 presents a detailed breakdown of these classifications.
+  Utterance Type Engagement Level Number of Utterances
+Explain Own Idea (EXP) 497
+Engage Others Idea (EOI) Total 696
+High Engagement 103
+Medium Engagement 205
+Low Engagement 388
+Uncorrelated 3,040
+Total Utterances 4,233
+Table 1: Classification Results by Utterance Type and Engagement Level
+  Name Utterances Label Engagement Level
+  Kimberly I thought you did just it by 5 because like (points to N’s notebook). engage others idea low
+Natalie Then I got to 5s and made 10, so I put 5 and 5 to (inaudible). And then I counted how much there was both and…it was 35. explain own idea -
+Kimberly Okay. Samantha go next. uncorrelated -
+Samantha What I did was right here.
+I did yesterday and then I didn’t have enough for everyone.
+And then I passed the last 10 because if I put it in with these, wouldn’t it be equal and then I split it.
+And then I carried it in and then I divided by once, and then I counted it and then I got 1, 2, 3, 4, 5 and this was 30, and then I got 35.
+explain own idea -
+Kimberly What I did was like I did like Samantha’s but I just labeled one day and two day, and then I just made a cross.
+Then, here I did the same thing, I got the seven 10s, but when I end up with 2, I just cross 2 off and then the 10.
+But then I noticed this was like that so I turned them into little 10 ones.
+And then I got 1, 2, 3, 4, 5, 6, 7, 8, 9.
+10.
+I got 35.
+explain own idea -
+Natalie (has a look of disbelief) How do you know it was 35? engage others idea medium
+Kimberly Like I counted it. 10, 20. I already knew this was 20, 30 explain own idea -
+Natalie 40. uncorrelated -
+Kimberly (laughs)…30, 31, 32, 33, 34, 35. And that’s my answer for how I got to 35. (adjusts camera) And now can I go first for the second number share? explain own idea -
+Natalie Yes. uncorrelated -
+Table 2: Examples: Identifying labels (explain vs. engage).
 […]
-Figure 1: System and User Prompts for Utterance Classification.
-Includes guidelines, category definitions, and example dialogues for analyzing and classifying the contribution of utterances in collaborative problem-solving processes.
-Figure 1 illustrates the user and system prompts implemented in this study.
-The prompts define the utterance classification task, outline categorization rules, and provide illustrative dialogues to clarify distinctions between categories.
-The user prompt provides clear guidelines for classifying utterances into three main categories: ‘explain own ideas’, ‘engage others’ ideas’, or ‘uncorrelated’.
-For the ‘engage others’ ideas’ category, it further defines three levels of engagement (low, medium, high) and offers specific criteria for making classification decisions, including special cases and exceptions.
-The system prompt supports these guidelines by stressing the importance of accurate categorization, emphasizing subtle differences between labels, and ensuring consistency across classifications.
-To enhance output reliability and reduce variability in the model’s responses, we adjusted the temperature parameter to nearly zero, optimizing consistency in classification.
-2.3 Utterance Classification Results
-We conducted utterance classification using prompts carefully constructed, as illustrated in Figure 1.
-These prompts were designed incorporating comprehensive contextual information, including dialogue tasks, categorization rules, and classification guidelines.
-To optimize classification accuracy, we included five few-shot examples in the prompts, each structured to specify speaker identity, utterance content, classification label, engagement level, and relevant references.
-Complete prompt specifications, including the few-shot examples, are detailed in the Section C of the Supplementary Material.
-Instead of individual utterances, dialogue chunks segmented by timestamps from the original dataset were included as the inputs for classification within our prompts.
-This approach allows the prompts to capture conversational context effectively and inform the classification decisions.
-The classification process was executed on an NVIDIA RTX A6000 GPU with an Intel(R) Xeon(R) Gold 6426Y CPU in a Linux environment, requiring approximately 1.5 hours for completion.
-[…]
-  Utterance Type Engagement Level Number of Utterances
-Explain Own Idea (EXP) 1,335
-Engage Others Idea (EOI) Total 1,251
-High Engagement 107
-Medium Engagement 553
-Low Engagement 591
-Uncorrelated 1,544
-Total Utterances 4,130
-Table 1: Classification Results by Utterance Type and Engagement Level
-[…]
-  Name texts Label Engagement Level
-  Delilah Okay now. First I put 2 days to represent the days. And then I put 70 units in each group (gestures across the two groups on her notebook). explain own idea -
-Keila First I… uncorrelated -
-Delilah And I came up with 39 because each one had 39. explain own idea -
-Keila Did you put it one by one? engage others idea medium
-Delilah Yeah. explain own idea -
-Keila Mine got 35. engage others idea low
-Delilah So I got 1, 2, 3, 4, 5, 6, 7, 8, blah blah blah. (counting the marks in one group) explain own idea -
-Keila Did you put 1, 2, 3, 4, 5, 6? (gestures marking in alternating groups as she counts, demonstrating action in solving) engage others idea medium
-Delilah Yeah! And 10 divided by 2 is 3. I meant 39. explain own idea -
-Table 2: Examples: Identifying labels (explain vs. engage) and engagement levels.
-  Name texts Label Engagement Level
+  Name Utterances Label Engagement Level
   William Okay so first what I did with this is I did 70 and 2. So I had two days. 1 day, 2 days, and I counted all the way up to 70 by 2, and I counted. And so now I got 32. explain own idea -
 Julian I disagree with you cause I got 1, 2, 3, (…), 14 (counts his 14 rods) engage others idea high
-William Wait hold on. You don’t add 7 plus 7. If you add 7 plus 7 that would be 14. But are you trying to add? Or times? engage others idea high
+William Wait hold on. You don’t add 7 plus 7. If you add 7 plus 7 that would be 14. But are you trying to add? Or times? engage others idea medium
 Table 3: High-Level Engagement in Problem-Solving
-Second, Table 3 illustrates the system’s ability to capture complex interaction sequences.
-The example begins with a student’s explanation of their approach (classified as EXP), followed by a peer’s substantive disagreement (classified as high-level EOI).
-The subsequent probing question, which builds upon this disagreement, maintains the high-level EOI classification.
+Table 3 illustrates the ability of the system to capture complex interaction sequences.
+The example begins with a student’s explanation of their approach, which is classified as explain own idea.
+This is followed by a peer’s substantive disagreement, appropriately classified as engage with others’ ideas with high engagement level.
+The subsequent question of probing, which builds on this disagreement, is classified as Engage with others’ ideas with medium engagement.
 This sequence demonstrates the capacity of the system to identify sophisticated interactions where students progressively build upon and critically engage with each other’s mathematical reasoning.
-  Name texts Label Engagement Level
+Table 4 demonstrates the system’s capacity to distinguish between mathematical discourse and off-topic conversations within classroom settings.
+While off-topic remarks are classified as uncorrelated, mathematical explanations receive relevant classifications such as engage with others’ ideas.
+The example further illustrates how the system recognizes varying levels of engagement with peer contributions, identifying both medium- and low-level engagement instances.
+Subsequently, when students transition to explaining their own reasoning, the system correctly adjusts the classification accordingly.
+This sequence exemplifies the system’s ability to maintain classification accuracy across diverse types of classroom discourse while maintaining focus on mathematically relevant interactions.
+  Name Utterances Label Engagement Level
   Natalie Okay, let me go. uncorrelated -
 Kimberly I missed a number. uncorrelated -
-Natalie I’m looking at the hundreds chart. (refers to yellow sheet) And then I went to 75. explain own idea -
+Natalie I’m looking at the hundreds chart. (refers to yellow sheet) And then I went to 75. uncorrelated -
 Kimberly Why did you go to 75? engage others idea medium
@@
 Natalie …that was how much shells she found. explain own idea -
-Kimberly It was 70. engage others idea low
-Natalie On the second strategy. explain own idea -
+Kimberly It was 70. uncorrelated -
+Natalie On the second strategy. uncorrelated -
 Kimberly Oh! uncorrelated -
@@
 Samantha We are already done. uncorrelated -
-Kimberly We are at the strategies, but she and me realized that we strategies on the 200 chart. uncorrelated -
+Kimberly We are at the strategies, but she and me realized that we strategies on the 200 chart. explain own idea -
 Table 4: Differentiating Relevant and Off-Topic Remarks.
+2.4 Validation of Utterance Classification
+To evaluate the reliability and validity of our utterance classification framework, we implemented a comprehensive multi-faceted validation approach.
+This validation strategy includes three primary analytical components: inter-model consistency analysis, human annotation comparison, and uncertainty quantification analysis.
+The inter-model consistency analysis assesses agreement levels between different LLMs, while human annotation comparison examines the consistency between human judgments and LLM-generated classifications.
+Additionally, we conducted both inter-model and intra-model uncertainty analyses to systematically identify instances requiring further review and to examine individual model confidence levels.
+These complementary approaches provide a thorough assessment of classification quality.
+2.4.1 Inter-Model Consistency Analysis
+We conducted a comprehensive inter-model consistency analysis to evaluate the degree of agreement between LLMs in their classification decisions.
+This analysis employed two established agreement metrics that provide complementary perspectives on model alignment: Cohen’s $\kappa$ (Cohen, 1960) and Fleiss’ $\kappa$ (Fleiss, 1971).
+Cohen’s kappa
+We utilized Cohen’s $\kappa$ coefficient (Cohen, 1960) to assess pairwise agreement between individual model pairs, which enables detailed examination of specific model relationships and identification of which models demonstrate the strongest concordance in their classification patterns.
+Specifically, Cohen’s $\kappa$ coefficient $\kappa_{Cohen}$ quantifies the degree of agreement between two classification models while accounting for chance agreement.
+Cohen’s $\kappa$ coefficient is defined as:
+$\kappa_{Cohen}=\frac{p_{o}-p_{e}}{1-p_{e}}$
+where $p_{o}$ represents the observed proportional agreement between the two models under comparison, while $p_{e}$ denotes the expected proportional agreement that would occur due to chance alone.
+The metric provides a chance-corrected measure of agreement that operates within a standardized range from -1, indicating perfect disagreement, to +1, representing perfect agreement.
+A coefficient value of 0 means that the observed agreement is equivalent to what would be expected through random chance alone.
+The interpretation of Cohen’s $\kappa$ values follows established conventions as originally proposed by Landis and Koch (1977), with specific thresholds and their corresponding agreement levels detailed in Table 5.
+$\kappa$ Statistic Strength of Agreement
+<< 0.00 Poor
+0.00 - 0.20 Slight
+0.21 - 0.40 Fair
+0.41 - 0.60 Moderate
+0.61 - 0.80 Substantial
+0.81 - 1.00 Almost perfect
+Table 5: Standardized interpretation framework for Cohen’s $\kappa$ coefficient values, providing qualitative assessments of inter-rater agreement strength.
+Figure 1 presents a heatmap visualization of pairwise Cohen’s $\kappa$ coefficients computed across all possible model combinations in our evaluation set.
+The analysis demonstrates predominantly strong agreement levels throughout the model pairs, with the majority of coefficients achieving substantial to almost perfect agreement ranges ($\kappa_{Cohen}>0.61$).
+However, a notable pattern emerges in the performance of Llama-3.1, which consistently exhibits lower agreement scores compared to all other models in the ensemble.
+This systematic deviation indicates that Llama-3.1 employs fundamentally different classification criteria or response generation mechanisms compared to the remaining six models.
+Figure 1: Heatmap visualization displaying pairwise Cohen’s $\kappa$ coefficients computed across all possible combinations of the seven LLMs included in our ensemble framework.
+Each cell quantifies the chance-corrected agreement between two specific models, with color intensity representing the magnitude of agreement strength.
+Darker color intensities correspond to higher agreement levels, while lighter shades indicate lower consensus between model pairs.
+Fleiss’ kappa
+Additionally, we employed Fleiss’ $\kappa$ (Fleiss, 1971) to evaluate multi-rater agreement across all seven models simultaneously, which yields a measure of overall ensemble consistency.
+Specifically, Fleiss’ $\kappa$ extends Cohen’s $\kappa$ framework to accommodate multiple raters in a single comprehensive analysis. Fleiss’ $\kappa$ is defined as:
+$\kappa_{Fleiss}=\frac{\bar{P}-\bar{P_{e}}}{1-\bar{P_{e}}}$ (1)
+where $\bar{P}$ represents the mean observed agreement across all classification items, while $\bar{P_{e}}$ denotes the mean expected agreement that would occur under the statistical assumption of independence among all seven models.
+This metric captures ensemble-wide consistency patterns that cannot be detected through pairwise comparisons alone, providing critical insights into the overall coherence.
+Group Fleiss’ $\kappa$
+LLMs (All) 0.6575
+Open-source models 0.6033
+Commercial API models 0.7354
+Table 6: Fleiss’ $\kappa$ coefficients demonstrating multi-rater agreement levels across different model grouping.
+The comprehensive analysis includes three distinct categories: the complete seven-model ensemble represented by ”LLMs (All),” the subset of open-source models (“Open-source Models”), and the collection of commercial API models (“Commercial API Models”).
+All calculated coefficients achieve substantial agreement levels within the established range of 0.61-0.80, providing empirical evidence of reliable consensus across diverse model architectures and confirming the robustness of our classification methodology regardless of model type or commercial availability.
+This comprehensive inter-model consistency analysis demonstrates robust agreement across our evaluation framework, with all models achieving substantial levels of consensus that validate the reliability of our classification approach.
+Fleiss’s kappa values consistently exceeded 0.60 across different model groupings, indicating strong agreement that surpasses conventional thresholds for acceptable inter-rater reliability.
+In particular, commercial API models demonstrated superior consensus ($\kappa_{Fleiss}$ = 0.7354) compared to their open source counterparts ($\kappa_{Fleiss}$ = 0.6033), suggesting that standardized development practices and potentially shared training methodologies within commercial ecosystems contribute to more aligned classification behaviors.
+The consistent identification of Llama-3.1 as an outlier in pairwise comparisons is a significant finding that merits further investigation of its underlying architectural or training characteristics that differentiate its classification approach from the majority of the ensemble.
+These results provide strong empirical support for the validity of our multi-model evaluation framework and establish confidence in the robustness and generalizability of our experimental findings across diverse model architectures and development paradigms.
+2.4.2 Human-LLM Classification Comparison
+To validate the reliability of our automated classification framework, we additionally applied a human annotation assessment and examined the consistency between human judgments and LLM-generated results.
+Specifically, we created standardized labeling guidelines that align precisely with the classification criteria used in our LLM prompts, ensuring methodological consistency between the automated and manual evaluation processes.
+Eight human annotators, comprising seven graduate students and one undergraduate student, participated in the validation study using a standardized web-based annotation platform specifically designed to minimize inter-annotator variability and maintain consistent labeling conditions.
+The human annotation dataset comprised six strategically selected dialogue episodes that maximized the representation of the participants in classroom interactions.
+This focused subset was selected due to time and resource constraints, as comprehensive annotation of the complete dataset collected during the five-month instructional period would have exceeded the available capacity.
+To establish quantitative measures of annotation reliability, we implemented both Cohen’s $\kappa$ and Fleiss’ $\kappa$ coefficients for systematic assessment of inter-annotator agreement patterns.
+The heatmap visualization displaying pairwise Cohen’s $\kappa$ values between human annotators and LLMs is presented in Section LABEL:supp-sec:uncertainty of the Supplementary Material.
+Group Fleiss’ $\kappa$
+Human Annotators 0.3509
+LLMs (All) 0.5070
+Commercial API Models 0.5933
+Table 7: Fleiss’ $\kappa$ Coefficients for human annotator agreement.
+Results are based on the same 5 selected blocks (out of 33 total blocks) used for human coding, while Table 6 uses all utterances across the full dataset.
+Table 7 presents the Fleiss’ $\kappa$ coefficients across distinct annotator groups, revealing significant variations in classification consistency.
+The analysis demonstrates that human annotators achieved the lowest inter-rater agreement ($\kappa_{Fleiss}$ = 0.3509), indicating substantial disagreement in their classification decisions.
+In contrast, the complete LLM ensemble exhibited moderate consistency levels ($\kappa_{Fleiss}$ = 0.5070), while commercial API models demonstrated the highest agreement ($\kappa_{Fleiss}$ = 0.5933) among all the evaluated groups.
+These findings illuminate the inherent challenges associated with dialogue classification tasks, where subjective interpretation and contextual ambiguity contribute to variability in human judgment.
+The observed pattern suggests that multiple valid interpretations may exist for individual utterances, reflecting the complex nature of conversational analysis rather than deficiencies in human annotation capabilities.
+The higher consistency demonstrated by automated classification systems, particularly commercial models, indicates that standardized algorithmic approaches may provide more reproducible labeling results compared to human annotators.
+2.4.3 Uncertainty Quantification
+Inter-Model Uncertainty Quantification
+To systematically quantify the degree of disagreement among models and identify utterances requiring expert review, we implemented Shannon entropy (Shannon, 1948) as our primary uncertainty measure.
+This information-theoretic approach enables consistent measurement of classification uncertainty by analyzing the probability distribution of label choices across multiple LLM outputs, where each label’s probability is calculated as the proportion of models selecting that label.
+For a given utterance with classification probabilities $\{p_{1},p_{2},\dots,p_{n}\}$ distributed across $n$ possible classification labels, the Shannon entropy, $H(X)$, is defined as:
+$H(X)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i},$ (2)
+where the probability $p_{i}$ for each label category is empirically estimated through ensemble voting as:
 […]
-2.4 Validation of Utterance Classification
-For the validation of our prompt engineering approach, we focused on assessing the alignment between intended classifications and actual outcomes.
-Due to the absence of pre-labeled data in our classroom dialogue dataset, we utilized the DailyDialog dataset (Li et al., 2017) as our external validation source.
-This dataset provides a structured framework for dialogue act classification, categorizing utterances into four distinct types: inform, questions, directives, and commissive.
-To evaluate our methodology, we developed dialogue act classification (DAC) specific prompts for this dataset, incorporating five dialogue chunks examples to create few-shot examples prompts.
-We compared the performance of these few-shot prompts against zero-shot approaches, enabling us to quantify the impact of including dialogue examples in the classification process.
-To establish the effectiveness and generalizability of our prompt engineering approach, we conducted comprehensive comparisons with established deep learning models.
-Context-Aware Self-Attention model (Raheja and Tetreault, 2019, CASA) employs a self-attention mechanism integrated with hierarchical deep neural networks to recognize dialogue acts within conversational contexts.
-The Deep Co-Interactive Relation Network (Qin et al., 2020, DCR-Net) employs a hierarchical encoder and relation layer to integrate contextual and interaction information for dialogue act classification.
-A co-interactive graph attention network model (Qin et al., 2021, Co-GAT) jointly performs DAC and sentiment classification, leveraging both contextual and interaction information through cross-utterance and cross-task connections.
-Additionally, Zhao et al.
-(2023) presented classification results using prompt engineering with the GPT-3.5-turbo model, and we compared our results against these presented results.
-We evaluated classification performance using three complementary metrics: macro-F1, weighted-F1 scores, and overall accuracy.
-Macro-F1 computes the F1 score for each class individually and then averages them, providing a balanced measure of performance across all classes, regardless of their frequencies.
-This is particularly useful in imbalanced datasets, where certain classes may dominate.
-On the other hand, Weighted-F1 calculates the F1 score for each class and takes into account the number of true instances for each class, giving more importance to classes with more instances.
-These metrics help assess the model’s performance from different perspectives, considering both class balance and overall accuracy.
-  Model Acc Macro-F1 Weighted-F1
-  CASA - 0.78 -
-DCR-Net + Co-Attention - 0.79 -
-Co-GAT - 0.79 -
-ChatGPT 1-shot + P.E 0.67 - 0.65
-ChatGPT 3-shot + P.E 0.73 - 0.71
-Llama-3.1-8B-Instruct Zero-shot 0.69 0.57 0.68
-Llama-3.1-8B-Instruct Few-shot 0.75 0.66 0.74
-Table 5: Comparison of Prompt Engineering results with other baselines on the DAC task.
-Experimental results for CASA, DCR-Net + Co-Attention, and Co-GAT are partly cited from Qin et al.
-(2021) and Zhao et al.
-(2023).
-P.E.
-stands for prompt engineering where the 1-shot and 3-shot settings use 1 and 3 examples in the prompt, respectively.
-Table 5 presents our comparative analysis of prompt engineering approaches and state-of-the-art deep learning models.
-Results demonstrate that the few-shot prompt engineering method, incorporating five dialogue chunks as examples, achieves substantially improved performance compared to the zero-shot approach.
-While some comparative studies did not report all evaluation metrics, we have calculated a comprehensive set of metrics for our prompt engineering results to facilitate thorough comparison across different models.
-This analytical approach ensures equitable performance assessment across varying methodologies.
-The comparative analysis reveals that prompt engineering performance, while effective, may not fully match that of deep learning models due to fundamental methodological distinctions.
-Deep learning approaches benefit from extensive optimization using pre-constructed labeled datasets and generate specialized embeddings tailored to specific tasks.
-In contrast, prompt engineering operates solely through label descriptions without prior training data, relying on the inherent language understanding capabilities of the model.
-This distinction becomes particularly significant in contexts where classification criteria exhibit variability due to multiple annotators or crowdsourcing processes.
-Nevertheless, these differences suggest that prompt engineering may be especially valuable in generating consistent labels, as it helps maintain classification criteria across datasets.
+The resulting entropy values operate within a defined range from 0, indicating perfect consensus among all models, to $\log_{2}(n)$, representing maximum uncertainty where models are equally divided across all possible labels.
+Utterances exhibiting high entropy values signal substantial inter-model disagreement and are systematically flagged as priority candidates for expert human annotation.
+This entropy-driven selection strategy enables efficient allocation of limited annotation resources while maintaining rigorous quality standards, effectively reducing the computational and human burden associated with comprehensive manual review processes.
+Figure 2: Distribution of Shannon entropy across seven LLMs for inter-model uncertainty quantification
+Figure 2 illustrates the distribution of Shannon entropy values derived from classification results across our seven LLMs.
+The analysis of 4,233 total utterances reveals that approximately 57 percent (2,433 utterances) achieved complete consensus among all models, demonstrating substantial inter-model agreement that validates the reliability of our classification framework.
+The calculated 95th percentile entropy threshold of 1.38 represents classification scenarios where four of the seven models converge on identical label assignments, indicating that the vast majority of utterances generate strong consensus patterns across the ensemble.
+Conversely, the top 5 percent of utterances exhibiting the highest entropy values reveal significant inter-model disagreement, creating natural breakpoints where either appropriate post-processing techniques or expert human annotation becomes essential for resolving classification ambiguity.
+In this study, we employed majority voting among commercial LLM APIs as a post-processing approach, given their higher consistency compared to other models.
+Intra-Model Uncertainty Quantification
+To evaluate model confidence with greater precision, we implemented an intra-model uncertainty assessment with two distinct approaches.
+The first approach employed sequence-level probability analysis, which measures the probability of complete classification labels by computing normalized logit scores across all possible label sequences (e.g., “Explain Own Idea,” “Engage with Others’ Ideas” with varying engagement levels, or “Uncorrelated”) and then quantifies the overall confidence in the final classification decision.
+The second approach utilized token-level probability analysis, which examines the probability distribution at each token generation step throughout the classification process, effectively replicates the internal decision-making mechanism of the LLM and provides detailed insights into specific decision points where uncertainty emerges.
+We applied Shannon entropy in Equation 2 as our primary uncertainty quantification metric for both analytical approaches.
+These complementary analytical approaches illuminate the distinct dimensions of model uncertainty, with sequence-level entropy that measures holistic linguistic confidence in complete label generation, and token-level entropy that identifies specific decision points where models demonstrate hesitation.
+Comprehensive sequence-level and token-level analysis results, model comparisons, and detailed methodological specifications are available in Section LABEL:supp-sec:uncertainty of the Supplementary Material.
+2.4.4 Summary
+Our comprehensive validation framework demonstrates the robustness and reliability of the LLM-based utterance classification approach through multiple analytical perspectives.
+Inter-model consistency analysis revealed substantial agreement levels across the ensemble, with commercial API models achieving the highest consensus ($\kappa_{Fleiss}$ = 0.7354) among all evaluated LLMs.
+Inter-model uncertainty quantification using Shannon entropy demonstrated that 57% of utterances achieved complete consensus across all models, while the top 5% with high uncertainty were effectively addressed through majority votes among commercial models.
+Intra-model uncertainty analysis provided complementary insights into model confidence patterns, revealing distinct uncertainty profiles across different model architectures.
+Comparative analysis with human annotations yielded particularly significant insights regarding classification consistency patterns.
+LLM ensembles achieved markedly higher inter-annotator agreement ($\kappa_{Fleiss}$ = 0.5070-0.5933) compared to human annotators ($\kappa_{Fleiss}$ = 0.3509), highlighting the substantial variability inherent in human interpretation of dialogue data.
+This finding demonstrates the considerable value of automated classification approaches in addressing the inherent subjectivity challenges that characterize dialogue annotation tasks, where human judgments frequently vary due to individual interpretative differences and contextual perspectives, thus supporting the rationale for LLM-based classification in reproducible interaction analysis.
+These validation results collectively establish strong confidence in our classification framework’s ability to generate reliable and methodologically sound labels for subsequent network analysis applications.
 3 Network Analysis of Classroom Interaction
@@
 We formalized student interactions using an adjacency matrix, where the matrix elements indicate the frequency of specific interaction types between two students during the observation period.
-We then visualized the constructed network structure by using nodes to represent individual students and directed weighted edges between nodes to represent the nature and frequency of the interactions between the two students.
+We then visualized the constructed network structure using nodes to represent individual students and directed weighted edges between nodes to represent the nature and frequency of the interactions between the two students.
 For instance, when a student offers multiple explanations to a peer, a directed edge is formed from the explaining student (source) to the listening student (target), with the edge weight reflecting the total number of explanatory interactions.
-Engagement interactions similarly generate directed edges from the engaging student to the idea originator, with edge weights reflecting interaction frequency.
+Engagement interactions similarly generate directed edges from the engaging student to the idea originator, with edge weights reflecting the interaction frequency.
 To capture the varying intensity of student engagement more precisely, we implemented a differential weighting scheme for these interactions.
 While explanatory interactions maintained equal weights, we applied multiplication factors to different levels of engagement: medium-level engagements received double weight, and high-level engagements received triple weight relative to low-level engagements.
-This weighted approach ensures that more substantive forms of engagement exert proportionally greater influence in the network structure, better reflecting their educational significance.
+This weighted approach ensures that more substantive forms of participation exert proportionally greater influence on the structure of the network, better reflecting their educational importance.
 While this network construction process allows us to represent valuable aspects of classroom interactions, analyzing classroom conversation data in a network also presents unique analytical challenges.
@@
 3.1.1 Network Graphs
-Figure 2 illustrates these network transformations, where each node represents a student and edges represent their interactions.
-Node colors indicate students’ standardized mathematics performance, utilizing a spectrum from blue (lower scores) to red (higher scores).
-Edge thickness corresponds to interaction frequency between student pairs.
+Figure 3 illustrates these network transformations, where each node represents a student, and the edges represent their interactions.
+The node colors indicate students’ standardized mathematics performance, utilizing a spectrum from blue (lower scores) to red (higher scores).
+The edge thickness corresponds to the interaction frequency between pairs of students.
 (a) Explain Own Idea Network, Pre-test
@@
 We then examined network centrality, using metrics designed for directed weighted networks, accounting for both directional flow and interaction intensity.
-These include PageRank (Page et al., 1999), hub-authority scores (Kleinberg, 1999), in-degree and out-degree, betweenness (Freeman, 1977), closeness (Bavelas, 1948), and eigenvalue centrality (Bonacich, 1972).
+These include PageRank (Page et al., 1999), hub-authority scores (Kleinberg, 1999), in-degree and out-degree, betweenness (Freeman, 1977), closeness (Bavelas, 1948), and eigenvector centrality (Bonacich, 1972).
 Each centrality measure captures distinct aspects of student engagement in the network.
@@
 In-degree and out-degree centrality measure the weighted volume of received and initiated interactions respectively, identifying key participants in collaborative discourse.
-Eigenvalue centrality assesses each student’s influence by considering the influence levels of their connected peers.
+Eigenvector centrality assesses each student’s influence by considering the influence levels of their connected peers.
 Betweenness and closeness centrality measures require a different approach to weight calculation, as they interpret edge weights as distances rather than strengths.
@@
 Pagerank Indegree Outdegree Betweenness Closeness Eigen Hub Authority
-  Angelina 0.1033 57 61 32 0.3695 0.2245 0.0378 0.0654
-Armando 0.0414 46 46 34 0.4829 0.2242 0.1974 0.0198
-Daniel 0.0286 48 37 27 0.4728 0.1811 0.0216 0.6029
-Diego 0.0451 75 47 115 0.5616 0.5980 0.0147 0.7355
-Esther 0.0479 92 105 76 0.5176 0.8422 0.0556 0.6152
-Gio 0.0515 56 139 145 0.5454 0.2639 1.0000 0.0171
-Isabel 0.0330 42 40 29 0.4775 0.2533 0.2659 0.0570
-Jessica 0.0522 71 26 57 0.4916 0.7276 0.0133 0.5414
-Julian 0.0530 87 51 52 0.5046 0.3090 0.0385 0.9111
-Keila 0.0229 35 22 0 0.3728 0.1329 0.0080 0.0877
-Kevin 0.0484 89 55 71 0.4571 0.2615 0.0045 1.0000
-Kimberly 0.0481 51 93 51 0.4613 0.2644 0.1562 0.0420
-Marco 0.0518 49 73 18 0.3276 0.1015 0.0067 0.2131
-Maria 0.0607 41 53 61 0.4259 0.2476 0.1405 0.0570
-Melissa 0.0184 25 46 0 0.3872 0.2847 0.1741 0.0217
-Mya 0.0665 92 135 142 0.5677 1.0000 0.6806 0.0618
-Natalie 0.0343 60 37 4 0.4005 0.4354 0.0421 0.2712
-Ruben 0.0481 59 74 43 0.4086 0.1663 0.5060 0.0051
-Samantha 0.0965 77 38 5 0.3629 0.2925 0.0247 0.0741
-William 0.0485 62 36 7 0.3242 0.1290 0.1091 0.0133
-Table 6: Centrality measures for the EXP (Explain Own Idea) network.
+Angelina 0.0857 28 31 11 0.1725 0.2081 0.3438 0.1159
+Armando 0.0223 8 8 0 0.1283 0.0694 0.0410 0.0258
+Daniel 0.0229 10 8 0 0.1295 0.0129 0.0077 0.0199
+Diego 0.0448 23 23 69 0.2064 0.4074 0.1418 0.4323
+Esther 0.0730 54 38 46 0.2377 0.7694 0.3451 1.0000
+Gio 0.0474 18 28 79 0.1828 0.0434 0.0718 0.0218
+Isabel 0.0388 24 28 109 0.2384 0.0865 0.2362 0.2082
+Jessica 0.0669 35 16 27 0.2047 0.6714 0.0684 0.7269
+Julian 0.0513 28 28 115 0.2130 0.0430 0.0516 0.1074
+Keila 0.0173 12 18 0 0.1916 0.0571 0.0765 0.1979
+Kevin 0.0307 18 9 10 0.1327 0.0092 0.0030 0.0169
+Kimberly 0.0517 38 54 123 0.2369 0.1693 0.5539 0.1953
+Marco 0.0425 19 28 27 0.1569 0.0072 0.0134 0.0122
+Maria 0.0548 15 23 7 0.1710 0.2063 0.2442 0.0445
+Melissa 0.0198 10 15 0 0.1560 0.1180 0.2284 0.1004
+Mya 0.1094 57 63 117 0.2340 1.0000 1.0000 0.4497
+Natalie 0.0478 30 36 57 0.1870 0.3322 0.3351 0.4367
+Ruben 0.0252 10 18 4 0.1243 0.0022 0.0059 0.0031
+Samantha 0.0928 47 25 36 0.1845 0.2919 0.0892 0.5415
+William 0.0548 34 21 55 0.1752 0.0200 0.0308 0.0284
+Table 8: Centrality measures for the EXP (Explain Own Idea) network. The top three values in each column are highlighted in bold.
 Pagerank Indegree Outdegree Betweenness Closeness Eigen Hub Authority
-  Angelina 0.0307 48 48 36 0.2968 0.0045 0.0006 0.0030
-Armando 0.0431 96 50 9 0.6094 0.4061 0.0796 0.0262
-Daniel 0.0358 36 33 0 0.4462 0.1607 0.0215 0.0174
-Diego 0.0556 96 128 135 0.6766 0.5410 0.0292 0.4539
-Esther 0.0964 192 75 110 0.6310 1.0000 0.0162 1.0000
-Gio 0.0744 134 75 125 0.5837 0.3982 0.0477 0.0443
-Isabel 0.0377 73 38 17 0.5011 0.2453 0.0999 0.0430
-Jessica 0.0278 59 80 90 0.5569 0.1406 0.0098 0.1212
-Julian 0.0600 71 152 57 0.5309 0.2287 0.0521 0.0290
-Keila 0.0228 33 40 0 0.4228 0.1101 0.0055 0.0623
-Kevin 0.0615 75 86 38 0.4894 0.2484 0.0398 0.0203
-Kimberly 0.0537 73 97 63 0.5287 0.2807 0.2482 0.0173
-Marco 0.0707 104 54 10 0.4116 0.2906 0.0092 0.0144
-Maria 0.0275 51 77 68 0.5223 0.0788 0.0449 0.0060
-Melissa 0.0481 59 58 11 0.5133 0.3470 0.3616 0.0274
-Mya 0.0646 119 175 167 0.6586 0.6136 1.0000 0.0229
-Natalie 0.0329 47 67 0 0.4311 0.1986 0.0909 0.1434
-Ruben 0.0629 82 76 29 0.4609 0.2486 0.0209 0.0447
-Samantha 0.0276 37 47 0 0.2647 0.0019 0.0044 0.0002
-William 0.0662 71 100 20 0.4313 0.2460 0.0153 0.0144
-Table 7: Centrality measures for the EOI (Engage Others’ Idea) network.
-Tables 6 and 7 illustrate how different centrality measures reveal distinct patterns of student engagement in the interaction networks.
-For example, in the explanation network, Mya demonstrates high PageRank centrality, indicating her frequent interactions with influential peers.
-Esther’s elevated in-degree and out-degree scores reflect her active participation as both recipient and initiator of mathematical explanations.
-Notably, Gio’s high betweenness centrality suggests a crucial role in facilitating information flow between different student groups, effectively bridging distinct classroom communities.
-Analysis of Table 7 reveals distinct interaction patterns within the Engagement network, with several students demonstrating notable centrality profiles.
-Esther’s consistently high scores across centrality metrics indicate her comprehensive engagement in mathematical discourse, both influencing and responding to peers’ ideas.
-Gio maintains his bridging role while actively participating in peer discussions, while Mya demonstrates strong bilateral engagement through both contributing to and building upon others’ mathematical reasoning.
-Additionally, Armando, Diego, and Julian’s elevated eigenvector centrality and hub scores suggest their significant influence within localized peer groups, indicating the presence of distinct collaborative clusters within the classroom.
-Our analysis shows that high network centrality does not necessarily correlate with improved mathematical performance.
-Mya’s case illustrates this complexity: despite her prominent position in both interaction networks, her mathematical achievement did not improve, possibly due to her frequent interactions with lower-performing peers.
-This suggests that the relationship between network centrality and academic performance is not straightforward and the specific composition of local interactions - such as which students interact with whom - appears to play a more significant role.
-The influence of classroom interactions on learning outcomes appears to depend more on the composition of immediate peer groups than on students’ overall network prominence.
-This observation underscores the importance of examining both the quality and context of student interactions within local network structures, rather than relying exclusively on global centrality metrics.
+Angelina 0.0354 42 38 36 0.1892 0.0020 0.0001 0.0054
+Armando 0.0456 53 26 41 0.3819 0.1221 0.0618 0.0190
+Daniel 0.0450 27 23 35 0.3289 0.1590 0.0026 0.0290
+Diego 0.0412 68 68 126 0.4468 0.2104 0.0075 0.9682
+Esther 0.0508 80 26 31 0.3907 0.1904 0.0063 1.0000
+Gio 0.0878 82 63 148 0.4118 0.7956 0.0649 0.0100
+Isabel 0.0554 48 22 17 0.3347 0.6473 0.0433 0.0228
+Jessica 0.0225 35 32 69 0.3278 0.0437 0.0010 0.3052
+Julian 0.1053 81 92 93 0.3811 1.0000 0.0109 0.0650
+Keila 0.0400 38 35 0 0.2372 0.0445 0.0018 0.0723
+Kevin 0.0454 37 44 30 0.2873 0.2871 0.0014 0.0174
+Kimberly 0.0641 54 63 38 0.2682 0.0651 0.1144 0.0060
+Marco 0.0585 69 44 5 0.2965 0.6780 0.0012 0.0251
+Maria 0.0261 27 41 70 0.3044 0.0157 0.0497 0.0014
+Melissa 0.0390 28 39 1 0.2515 0.0423 0.1496 0.0079
+Mya 0.0507 54 139 167 0.4264 0.1348 1.0000 0.0064
+Natalie 0.0472 51 41 47 0.2745 0.0613 0.0193 0.2643
+Ruben 0.0462 41 42 3 0.2588 0.2731 0.0074 0.0010
+Samantha 0.0313 30 39 0 0.1733 0.0010 0.0018 0.0001
+William 0.0623 51 79 34 0.3228 0.6860 0.0263 0.0040
+Table 9: Centrality measures for the EOI (Engage Others’ Idea) network. The top three values in each column are highlighted in bold.
+Tables 8 and 9 demonstrate how multiple centrality measures provide complementary insights into different patterns of student engagement within classroom interaction networks.
+The analysis reveals the differentiated roles that individual students take in facilitating mathematical discourse.
+Within the explanation network, Mya emerges as the most influential contributor, achieving the highest scores across the PageRank, out-degree, eigenvector, and hub centrality measures.
+On the other hand, Esther occupies a distinct but equally important position, ranking highest in in-degree and authority centrality scores.
+This pattern suggests that she serves as a focal point for peer explanations, with many students directing their mathematical reasoning toward her for consideration or validation.
+Kimberly demonstrates yet another critical network function, exhibiting the highest betweenness centrality score among all participants.
+This metric highlights her essential role as a connector who bridges other separate groups of students, thereby facilitating information flow and maintaining network cohesion throughout the classroom community.
+Table 9 shows distinct interaction patterns in the engagement network that highlight the diverse roles of students in peer-to-peer mathematical discourse.
+Mya again emerges as the primary engagement initiator, achieving the highest scores in out-degree, betweenness, and hub centrality measures.
+These convergent metrics demonstrate her exceptional capacity for outgoing engagement and her pivotal position in the dissemination of ideas throughout the classroom network.
+Gio complements this pattern by demonstrating the highest PageRank score alongside top-tier performance in betweenness and degree centrality measures, establishing his critical function as a connector who bridges distinct subgroups within the classroom community.
+Esther ranks highest in in-degree and authority centrality while exhibiting comparatively lower out-degree and eigenvector scores.
+This profile indicates that she serves as a focal point for peer engagement, frequently receiving directed interactions from classmates despite maintaining a more selective approach to outward influence.
+Julian demonstrates influence through strategic positioning, achieving the top eigenvector centrality score alongside high-degree measures that suggest that his impact operates primarily through well-connected peer relationships rather than direct engagement volume.
+Melissa and Kimberly both show elevated hub scores that reflect their active participation in initiating discussions and amplifying peer contributions throughout the network.
+Our exploratory network analysis suggests that high network centrality does not necessarily translate into improved mathematical performance, indicating that prominent positioning within classroom interaction networks alone may be insufficient to guarantee learning outcomes.
+Mya’s case exemplifies this complex dynamic.
+Despite maintaining prominent positions across explanation and engagement networks, her mathematical achievement remained stagnant throughout the study period.
+This counterintuitive result may be attributable to the composition of her peer interactions, particularly her frequent engagement with lower-performing students who may not have provided the mathematical scaffolding necessary for academic advancement.
+These findings demonstrate that the relationship between network position and academic success operates through more nuanced mechanisms than simple connectivity metrics suggest.
+The quality and academic caliber of immediate peer connections appear to exert greater influence on learning outcomes than overall network prominence or global centrality measures.
+This discovery underscores the critical importance of examining local network structures and interaction patterns rather than relying exclusively on aggregate centrality indicators.
+It is important to stress that the above interpretations, though, are exploratory solely based on network centrality measures, and do not necessarily imply causal relationships.
 3.2 Network Mediation Model
 Following the descriptive analysis of classroom interaction networks through visualization and centrality metrics, we now proceed to a more rigorous examination of how these network structures influence student outcomes.
-While network centrality measures offer valuable insights into the roles and prominence of individual students within the interaction network, they do not directly reveal the mechanisms through which these interaction patterns impact learning.
+While network centrality measures offer valuable insight into the roles and prominence of individual students within the interaction network, they do not directly reveal the mechanisms through which these interaction patterns impact learning.
 To understand these mechanisms, we employ network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022).
@@
 Our study examines how classroom interaction networks mediate the influence of gender on mathematics achievement.
-Prior research has consistently investigated gender differences in mathematics performance, revealing complex patterns and influences (Lindberg et al., 2010; Else-Quest et al., 2010; Ghasemi and Burley, 2019), while Lindberg et al.
+Previous research has consistently investigated gender differences in mathematics performance, revealing complex patterns and influences (Lindberg et al., 2010; Else-Quest et al., 2010; Ghasemi and Burley, 2019), while Lindberg et al.
 (2010) demonstrate that overall gender gaps in mathematics performance are relatively small.
@@
 Else-Quest et al.
-(2010) emphasizes the role of societal context in gender disparities, noting smaller performance differences in countries with greater gender equality.
+(2010) emphasizes the role of the social context in gender disparities, noting smaller performance differences in countries with greater gender equality.
 These findings collectively suggest that gender differences in mathematics achievement stem from social rather than purely biological factors, underscoring the importance of analyzing classroom social networks and their mediating effects on academic performance.
 Based on this perspective, we hypothesize that the gender gap in mathematics performance ($X$ on $Y$) can be differentiated based on how individual students interact with others during mathematics discussion classes.
-Specifically, we apply network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022), using students’ interaction structures as a mediator ($M$) for gender ($X$) - math performance ($Y$) relationships.
-By incorporating the estimated latent positions ($z_{i}$, $w_{j}$) as mediators, our goal is to investigate the indirect effects of gender on mathematics performance through the network structure.
-Pre-test Scores as a Confounder
-In the current context, pre-test scores could serve as a useful confounder for the intended mediation analysis.
-However, the pre-test (California Standardized test; CST) and post-test (a test developed by Webb et al.
-(2023)) do not measure the same domains, while the assessment methods and criteria are also different between the two tests.
-Thus, the two test scores may not be entirely comparable (Liu et al., 2017).
-Figure 4: Scatter plot showing the relationship between pre-test (CST) and post-test scores.
-Further, there is an additional technical challenge: the pre-test scores do not seem to have a linear relationship with the post-test scores.
-As shown in Figure 4, the relationship is not clearly linear around the Basic (300-350) and Proficient (350-400) ranges of the pre-test.
-This means that higher scores in the pre-test may not predict higher post-test scores consistently, suggesting pre-test scores may not be a reliable predictor of the post-test scores.
-To account for students’ prior mathematical proficiency in our analysis, we constructed a binary proxy variable based on California Standards Test (CST) performance.
-Students who achieved scores of 350 or above — the state-designated threshold for proficiency [^2] https://www.starsamplequestions.org/about.html — were categorized as high-performing.
-This dichotomization aligns with state educational standards and provides a meaningful distinction in student achievement levels.
-We incorporate this proxy variable as a confounder in our mediation analysis to more precisely evaluate how network structures mediate gender-based differences in mathematics performance.
-Figure 5: Network mediation model path diagram.
-This model analyzes the effect of gender on math achievement mediated by network interactions.
-Latent positions of sender and receiver networks are used as mediators.
+Specifically, we apply network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022), using students’ interaction structures as mediators ($M$) for gender ($X$) - math performance ($Y$) relationships.
+By incorporating estimated latent positions ($z_{i}$, $w_{j}$) as mediators, our goal is to investigate the indirect effects of gender on mathematics performance through the network structure.
+To address potential confounding effects, we incorporated students’ prior mathematical proficiency as a confounding variable in our mediation analysis.
+We use California Standards Test (CST) performance, with students scoring 350 or above the state-designated proficiency threshold classified as high-performing and those below this threshold categorized as requiring additional support.
+This approach addresses non-linear relationships observed between pre-test and post-test performance measures while systematically controlling for baseline achievement differences that could otherwise confound the relationship between gender, network position, and learning outcomes.
+Detailed justification for our choice of the confounding variable is provided in Section LABEL:supp-sec:confounding of the Supplementary Material.
 Network Mediation Model
-We additionally consider predictor-mediator interaction effects to account for potential heterogeneity in mediator effects across varying predictor levels (VanderWeele and Vansteelandt, 2009; Valeri and VanderWeele, 2013).
-In this case, the indirect effects, mediated through the latent network structure, vary with gender ($x$).
+We also consider the effects of the predictor-mediator interaction to account for the potential heterogeneity in mediator effects at different predictor levels (VanderWeele and Vansteelandt, 2009; Valeri and VanderWeele, 2013).
+In this case, indirect effects, mediated through the latent network structure, vary by gender ($x$).
 By modeling this interaction, we account for the possibility that the indirect effect of gender on math performance is mediated differently depending on the students’ positions within the network.
-The resulting network mediation model can be expressed as:
+The resulting network mediation model can be expressed as follows:
 $\begin{split}E(M_{k}\mid x,c)&=a_{0k}+a_{1k}\,x+a_{2k}\,c,\\ E(Y\mid x,\mathbf{m},c)&=b_{0}+b_{1}\,x+\sum_{k}b_{2k}\,m_{k}+\sum_{k}b_{3k}\,x\,m_{k}+b_{4}\,c,\end{split}$ (4)
-where $x$ is the binary indicator for gender, $m_{k}$ represents the latent positions of students, and $c$ is the confounder indicating proficiency based on the CST score threshold.
-Specifically, the latent variables $m_{k}$ represent the coordinates of the latent positions for the network nodes: $m_{1}$ and $m_{2}$ correspond to sender position coordinates (x and y axes, respectively), while $m_{3}$ and $m_{4}$ represent receiver position coordinates (x and y axes, respectively).
+where $x$ is the binary indicator for gender, $m_{k}$ represents the latent positions of the students, and $c$ is the confounder indicating proficiency based on the CST score threshold.
+Specifically, the latent variables $m_{k}$ represent the coordinates of the latent positions for the network nodes: $m_{1}$ and $m_{2}$ correspond to the coordinates of the sender position (x and y axes, respectively), while $m_{3}$ and $m_{4}$ represent the coordinates of the receiver position (x and y axes, respectively).
 The model coefficients capture distinct aspects of the gender-performance relationship: $a_{1k}$ quantifies gender’s influence on network positioning, $b_{2k}$ measures how network positions affect mathematics performance, and $b_{3k}$ captures the predictor-mediator interaction between gender and network position, revealing how network-mediated effects vary by gender.
@@
 These components constitute the total effect (TE), providing a comprehensive measure of the impact of gender on math performance.
-The natural direct effect, the natural indirect effect, and the total effect, are given by:
+The natural direct effect, the natural indirect effect, and the total effect are given by:
 $\begin{split}NDE&=\left(b_{1}+\sum_{k}b_{3k}\left(a_{0k}+a_{1k}\,x^{*}+a_{2k}\,c\right)\right)(x-x^{*}),\\ NIE&=\left(\sum_{k}\left(b_{2k}\,a_{1k}+b_{3k}\,a_{1k}\,x\right)\right)(x-x^{*}),\\ TE&=NDE+NIE.\end{split}$ (5)
@@
 Our analysis examines how distinct types of classroom interactions mediate gender-based differences in mathematics performance by investigating two separate networks: the EXP network and the EOI network.
-For both networks, we incorporate students’ prior mathematics achievement as a confounder to account for pre-existing academic differences.
+For both networks, we incorporate students’ prior mathematics achievement as a confounder to account for preexisting academic differences.
 Through the application of our network mediation model to these distinct interaction networks, we aim to investigate whether the mediating effects of classroom network structures vary systematically based on interaction type.
@@
 To obtain the latent positions that serve as mediators, we employ the AMEN model (Hoff, 2021).
-The AMEN model is an extension of latent space model (LSM; Hoff et al., 2002), designed to capture more sophisticated network patterns, particularly in directed networks like classroom interactions.
+The AMEN model is an extension of the latent space model (LSM; Hoff et al., 2002), designed to capture more sophisticated network patterns, particularly in directed networks such as classroom interactions.
 This extension is particularly relevant for our analysis of classroom interactions, which are inherently directed.
 The AMEN model explicitly models directed relationships by providing distinct latent positions for each student as both a sender and a receiver.
-The AMEN model can be expressed as:
+The AMEN model can be expressed as
 $g(\mathbb{E}(y_{ij}))=\alpha_{i}+\beta_{j}+z_{i}^{T}w_{j},$ (6)
@@
 However, a Poisson model often suffers from overdispersion, where the observed variance exceeds the theoretical mean.
-The sender and receiver random effects can account for overdispersion to some degree by allowing the conditional mean to vary across participants (Gelman, 2007).
+Random effects of the sender and receiver can account for overdispersion to some degree by allowing the conditional mean to vary across participants (Gelman, 2007).
 However, severe overdispersion may not be fully addressable with the sender and receiver random effects.
-In our analysis of classroom interaction data, we identified a severe overdispersion issue, as detailed in Section D of the Supplementary Material.
+In our analysis of classroom interaction data, we identified a severe overdispersion issue, as detailed in Section LABEL:supp-sec:overdispersion of the Supplementary Material.
 To effectively address overdispersion, we propose using a negative binomial regression model, which introduces a dispersion parameter to explicitly account for the excess variability and allows the conditional variance to vary independently of the mean.
-By combining the negative binomial specification with sender and receiver random effects from the AMEM model, we can effectively capture both individual-level heterogeneity and the overdispersed nature of classroom interaction counts.
+By combining the negative binomial specification with sender and receiver random effects from the AMEN model, we can effectively capture both individual-level heterogeneity and the overdispersed nature of classroom interaction counts.
 The proposed AMEN model with a negative binomial regression structure for count data can be written as:
 $\log(\mu_{ij})=\alpha_{i}+\beta_{j}+z_{i}^{T}w_{j},$ (7)
-where $\mu_{ij}$ is the mean of the negative binomial distribution where $y_{ij}\sim\mbox{NB}(\mu_{ij},r)$ and $r$ of the negative binomial distribution controls the variance of the distribution, with higher values indicating greater dispersion relative to the mean.
-$\alpha_{i}$ and $\beta_{j}$ are sender and receiver random effects, respectively, and $z_{i}$ and $w_{j}$ are sender and receiver latent feature vectors, respectively.
+where $\mu_{ij}$ is the mean of the negative binomial distribution, where $y_{ij}\sim\mbox{NB}(\mu_{ij},r)$ and $r$ of the negative binomial distribution control the variance of the distribution, with higher values indicating a greater dispersion relative to the mean.
+$\alpha_{i}$ and $\beta_{j}$ are random effects of the sender and the receiver, respectively, and $z_{i}$ and $w_{j}$ are latent vectors of the characteristics of the sender and the receiver, respectively.
 3.4 Estimation Methods
@@
 AMEN Model Estimation
-In the first stage, we estimate the proposed AMEN model to embed the observed student interaction network into a latent space.
+In the first stage, we estimate the proposed AMEN model to embed the observed student-interaction network into a latent space.
 The prior distributions of the AMEN model parameters are specified as follows:
@@
 where $\sigma_{\alpha}^{2}$, $\sigma_{\beta}^{2}$, $\sigma_{z}^{2}$, $\sigma_{w}^{2}$, and $\sigma_{r}^{2}$ represent fixed prior variance terms for each corresponding parameter, and $I_{d}$ denotes the $d\times d$ identity matrix.
-In this study, we set $d=2$ for interpretability and visualization purposes.
-We draw the posterior samples of the AMEN model parameters and latent positions using a Metropolis-Hastings-within-Gibbs sampler.
-Additional details on the prior means, standard deviations, and the jumping rule for each parameter, as well as the starting points and overall convergence checks, are provided in the Section A of the Supplementary Material.
-Note that the inner product between latent vectors remains invariant under transformations such as rotation or reflection.
-To address this, we perform Procrustes matching (Gower, 1975) to align the latent vectors to a common reference position.
+We applied a data-driven approach to determine the latent space dimension $d$ for the AMEN model by conducting a systematic comparison of the model across candidate values $d\in\{2,3,4,5\}$ with the Bayesian Information Criterion (BIC), Deviance Information Criterion (DIC), and Watanabe-Akaike Information Criterion (WAIC) (see Table LABEL:supp-tab:dimension_selection in Section LABEL:supp-sec:amen of the Supplementary Material).
+We selected $d=2$ for the EXP (Explain Own Idea) network and $d=3$ for the EOI (Engage with Others’ Ideas) network, balancing goodness-of-fit and model complexity.
+Note that dimension selection in latent space modeling is an evolving area of research with no standard rules, and we therefore employed multiple information criteria to compare model performance across different dimensions.
+We draw posterior samples of the AMEN model parameters and latent positions using MCMC.
+Specifically, the Metropolis-adjusted Langevin algorithm (MALA) is applied to update the latent positions $z$ and $w$, while the remaining parameters are updated using a Metropolis-Hastings-within-Gibbs sampler.
+Additional details on the prior means, standard deviations, and the jumping rules for the model parameters, as well as the starting points and overall convergence checks, are provided in Section LABEL:supp-sec:amen of the Supplementary Material.
+Since the inner product between latent vectors remains invariant under transformations such as rotation or reflection, we performed Procrustes matching (Gower, 1975) to align the latent vectors to a common reference position.
 This alignment ensures that the latent vectors are consistently positioned, facilitating a clearer interpretation of their relationships and enhancing the validity of subsequent analyses.
@@
 In the second stage, we estimate a network mediation model.
-Specifically, the latent positions from the estimated AMEN model serve as mediators in the regression framework outlined in Equation (2).
+Specifically, the latent positions of the estimated AMEN model serve as mediators in the regression framework outlined in Equation (5).
 The network mediation model is estimated through Bayesian regression, with prior distributions specified as:
-[…]
-where $\Theta\in\{a_{01},\dots,a_{2k},b_{0},b_{1},b_{21}\dots,b_{3k},b_{4}\}$ represents all regression coefficients in Equation (1), $\epsilon_{i}$ represents the error term in the regression model, and $\sigma^{2}$ is the variance of error term.
+$\Theta\sim N(\bm{0},I),\quad\epsilon_{i}\sim N(0,\sigma^{2}),\quad\sigma^{2}\sim\text{Inverse-gamma}(0.001,0.001),$
+where $\Theta\in\{a_{01},\dots,a_{2k},b_{0},b_{1},b_{21}\dots,b_{3k},b_{4}\}$ represents all the regression coefficients in Equation (4), $\epsilon_{i}$ represents the error term in the regression model, and $\sigma^{2}$ is the variance of the error term.
 The network mediation model is estimated using the MCMCpack package (Martin et al., 2011) in R.
-The estimated regression coefficients are then used to calculate the natural direct effect (NDE) and natural indirect effect (NIE), as defined in Equation (2).
+The estimated regression coefficients are then used to calculate the natural direct effect (NDE) and the natural indirect effect (NIE), as defined in Equation (5).
 3.5 Network Analysis Results
@@
 Latent Position Structure
-Figures 6(a) and 6(b) illustrate the latent positions of students within both the EXP and EOI networks through two-dimensional visualizations.
+Figures 5(a) and 5(b) illustrate the latent positions of students within both the EXP and EOI networks through two-dimensional visualizations.
 In these representations, each student’s position is mapped as a point in two-dimensional space.
-The angle between student vectors indicates their interaction similarity - smaller angles represent greater similarity in interaction patterns.
-Additionally, vector magnitude represents each student’s relative influence within the network structure.
-Figure 2(a) reveals important bridging roles within the network structure, particularly through the positions of Gio and Isabel.
-Gio functions as a critical connector between two distinct student subgroups: one comprising Diego, Daniel, and Armando, and another consisting of Julian, William, and Marco.
-Similarly, Isabel facilitates connections between the subgroup of Melissa, Natalie, and Esther and the subgroup of Julian, William, and Marco.
-The latent vector visualization in Figure 6(a) captures these bridging relationships effectively, positioning Gio centrally between the connected subgroups while placing Isabel in closer proximity to the subgroup of Melissa, Natalie, and Esther.
-In Figure 2(c), Isabel serves a bridging role in connecting two distinct student clusters: the Keila and Kimberly group and the Marco and Julian group.
-In Figure 6(b), Isabel appears closer to the Keila, Kimberly, Melissa, Natalie, and Esther groups, reflecting the higher frequency of interactions within this group compared to the Marco and Julian cluster.
-The visualization also captures subtle relationship dynamics, such as Esther’s unidirectional engagement with William, which influences the latent space positioning of both Isabel and William through second-order interaction effects.
+The angle between the student vectors indicates their interaction similarity; smaller angles represent greater similarity in interaction patterns.
+Additionally, the magnitude of the vector represents each student’s relative influence within the network structure.
+The results suggest that Gio takes a particularly significant role as the primary intermediary linking two distinct clusters: the first comprising Diego and Armando, and the second encompassing Julian, William, and Marco.
+Similarly, Isabel performs a complementary bridging function by connecting the cluster of Keila, Kimberly and Esther with the Julian-William-Marco subgroup.
+The latent vector visualization presented in Figure 5(a) provides empirical support for these structural relationships, positioning Gio at a central location equidistant from both connected subgroups, while Isabel appears in closer proximity to the Keila-Kimberly-Esther cluster, indicating her primary affiliation while maintaining her bridging capacity to the adjacent subgroup.
+Figure 3(c) shows that Isabel serves as a bridge in connecting two distinct student groups: the Keila and Kimberly group and the Gio and Julian group.
+In Figure 5(b), Isabel is positioned between the Gio–William–Julian group and the Kimberly–Esther–Natalie cluster, indicating her intermediate role in the network.
+Daniel also serves as a bridge, linking the Kevin and Julian group and the Maria and Jessica group.
 These examples demonstrate the advantages of latent vector analysis in visualizing complex network data.
-By projecting interactions into a low-dimensional space, this approach uncovers hidden structures, including higher-order interactions and bridging relationships, while precisely representing both directionality and intensity of student connections.
-This methodology provides valuable insights into group dynamics, particularly for applications in educational research.
+By projecting interactions into a low-dimensional space, the present approach uncovers hidden structures, including higher-order interactions and bridging relationships, while precisely representing both directionality and intensity of student connections.
+Thereby, this methodology provides useful insights into group dynamics in education settings like the current example.
 3.5.2 Network Mediation Model Results
@@
 This analytical approach specifically evaluates how network structures mediate the relationship between gender and mathematics learning outcomes.
-The network mediation analysis quantifies the three key components: the natural indirect effect (NIE), the natural direct effect (NDE), and the total effect (TE).
-Tables 8 and 9 present the comprehensive results of these analyses.
+Network mediation analysis quantifies the three key components: NIE, NDE, and TE.
+To ensure robustness and reliability, we conducted the analysis using 10 independent MCMC chains and applied convergence diagnostics.
+In addition, to ensure consistent parameter interpretation across chains, all posterior samples were aligned with the chain exhibiting the lowest deviation information criterion (DIC) value, using the Maximum A Posteriori (MAP) estimate as the reference point for parameter alignment.
+This alignment procedure was completed before computing the final parameter estimates and credible intervals.
+Tables 10 and 11 present the comprehensive results of these analyses.
 Post.M Post.SD 2.5% 97.5%
-NIE 0.7758 1.2468 -1.7182 3.3649
-NDE -5.475 1.5948 -8.8699 -2.48 *
-TE -4.6992 0.8321 -6.4324 -3.056 *
+NIE 1.2317 (0.0324) 0.7538 (0.0107) -0.2966 (0.0306) 2.678 (0.0432)
+NDE -1.6922 (0.0319) 0.8183 (0.0107) -3.3055 (0.0442) -0.1052 (0.0308) *
+TE -0.4605 (0.0098) 0.3612 (0.0046) -1.187 (0.0119) 0.2604 (0.0131)
+Table 10: Network Mediation Model Results for EXP Network
+Analysis of the EXP network reveals a significant pattern, as shown in Table 10.
+NDE achieves statistical significance with a clear negative coefficient, indicating that female students experience lower mathematics performance through direct pathways independent of network-mediated influences.
+Although NIE does not reach conventional statistical significance thresholds, it demonstrates a notable positive coefficient, suggesting that gender influences explanation-seeking network positioning in ways that may benefit female students’ mathematical achievement.
+The findings suggest that explanation networks may function as collaborative learning structures where female students can leverage peer interactions to enhance their mathematical understanding, despite facing direct performance challenges.
+Post.M Post.SD 2.5% 97.5%
+NIE 0.5888 (0.0662) 0.9297 (0.0162) -1.2519 (0.0864) 2.4562 (0.0497)
+NDE -1.0528 (0.071) 1.1187 (0.0127) -3.1864 (0.0646) 1.1893 (0.0822)
+TE -0.464 (0.0059) 0.6794 (0.0111) -1.8654 (0.0173) 0.8884 (0.032)
 Table 11: Network Mediation Model Results for EOI Network
-Analysis of the EOI network, presented in Table 8, indicates that NDE is statistically significant and negative, while NIE through the network is not statistically significant.
-This negative NDE after controlling for proficiency of pre-test as a confounder, suggests that gender has a clear negative direct effect on post-test scores, meaning female students tend to achieve lower math scores independent of the mediating EOI network structure.
-Post.M Post.SD 2.5% 97.5%
-NIE 2.8653 1.3964 0.5089 5.9881 *
-NDE -6.9407 1.9286 -11.0408 -3.3803 *
-TE -4.0754 1.1155 -6.344 -1.7559 *
-Table 9: Network Mediation Model Results for EXP Network
-In contrast, for the EXP network, Table 9 reveals a different pattern.
-The negative NDE, which is consistent with the EOI network, shows that female students’ scores tend to be lower when considering gender independently of network effects.
-However, the positive NIE suggests that gender influences network structure in ways that enhance female students’ performance.
-These findings highlight the crucial role of network interactions, particularly within the EXP network, in mediating and potentially mitigating gender-based performance differences through collaborative learning processes.
+In contrast, analysis of the EOI network presented in Table 11 demonstrates that neither NDE nor NIE through network structure achieves statistical significance, as evidenced by 95% credible intervals that encompass zero.
+These results suggest that within the context of the EOI network, gender differences in mathematical achievement operate through alternative pathways that are not accounted for in our current network mediation framework.
+The results suggest that EOI interactions do not function as a significant mediating mechanism for gender-based performance differences.
+On the other hand, analysis of the EXP network reveals a contrasting pattern, as shown in Table 10.
+NDE achieves statistical significance with a clear negative coefficient, indicating that female students experience lower mathematics performance through direct pathways independent of network-mediated influences.
+Although NIE does not reach conventional statistical significance thresholds, it demonstrates a notable positive coefficient, suggesting that gender influences explanation-seeking network positioning in ways that may benefit female students’ mathematical achievement.
+The findings suggest that explanation networks may function as collaborative learning structures where female students can leverage peer interactions to enhance their mathematical understanding, despite facing direct performance challenges.
 4 Discussion
@@
 The prompt engineering approach enables automated classification of interaction patterns without requiring extensive pre-labeled datasets, significantly improving upon traditional manual labeling methods in both efficiency and scalability.
+Our extensive validation study supports these advantages, demonstrating that LLMs achieve high inter-model consistency even on complex dialogue classification tasks where human annotators show low agreement.
+Despite the inherent complexity of this task, our ensemble approach produces consistent labeling results while systematically identifying uncertain utterances through Shannon entropy-based uncertainty quantification.
 Building on this automated classification, our network analysis framework represents students as nodes and their interactions as edges, providing a comprehensive visualization and analysis of classroom information flow and engagement patterns.
-Through the projection of these networks onto latent spaces, we quantified structural characteristics that provide the understanding structure of classroom interactions.
-The network mediation analysis provides important insights into the relationship between classroom dynamics and student achievement.
-Our findings demonstrate that network mediation models possess unique capabilities for capturing intricate social patterns in educational environments which traditional analytical approaches often fail to identify due to the multidimensional nature of classroom interactions.
-This research opens several promising directions for future research, particularly in the incorporation of additional variables and the development of more sophisticated network representations, such as dynamic networks to capture temporal evolution of interactions or implement multilayer networks to represent diverse forms of student engagement simultaneously.
-Our methodological integration validates the effectiveness of combining prompt engineering with network analysis for understanding how student interactions influence learning outcomes.
+Through the projection of these networks onto latent spaces, we quantified the structural characteristics that provide the structure of the understanding of classroom interactions.
+Our contribution extends beyond traditional approaches through the integration of ensemble-based prompt engineering with network mediation analysis.
+To our knowledge, this combined approach with systematic uncertainty quantification and network-based structural analysis has not been applied in educational dialogue analysis, representing a novel methodological advancement that addresses both classification reliability challenges and the limitations of utterance-level analysis through relational pattern analysis.
+Network mediation analysis provides important information on the relationship between classroom dynamics and student achievement.
+Our findings demonstrate that network mediation models possess unique capabilities for capturing intricate social patterns in educational environments that traditional analytical approaches often fail to identify due to the multidimensional nature of classroom interactions.
+This research opens several promising directions for future research, particularly in the incorporation of additional variables and the development of more sophisticated network representations, such as dynamic networks to capture the temporal evolution of interactions or implement multilayer networks to represent diverse forms of student engagement simultaneously.
+Our methodological integration demonstrates the potential of combining prompt engineering with network analysis to understand how student interactions are associated with learning outcomes, with our mediation analysis being exploratory in nature and aiming to identify patterns and associations rather than establish causal relationships.
 Our findings suggest several promising directions for future research to deepen our understanding of classroom dynamics and refine analytical methodologies.
@@
 Teachers serve as essential facilitators of learning, significantly influencing classroom dynamics and student engagement patterns.
-Examining these teacher-student interactions would provide valuable insights into how instructional practices shape educational outcomes, particularly in the development of critical thinking skills and collaborative learning.
-Such analysis could inform the optimization of teaching strategies and contribute to enhanced classroom effectiveness.
-Future research could enhance the accuracy of interaction analysis by exploring model fine-tuning with high-quality labeled datasets, extending beyond our current prompt engineering approach.
-Such fine-tuning could significantly improve the model’s precision and generalizability in analyzing complex educational interactions.
-Additionally, the systematic integration of expert knowledge and carefully curated datasets into the prompt engineering process would strengthen the methodological reliability and validity.
-This enhanced approach could provide more nuanced insights into classroom interaction patterns while maintaining the efficiency advantages of automated analysis.
-A critical direction for future research involves developing automated prompt-generation methodologies.
-The current reliance on human judgment for prompt creation and selection introduces potential variability and bias into the analytical process.
-An automated approach that selects prompts based on their likelihood of generating desired outputs would enhance both the efficiency and consistency of the labeling process.
-Such automation would transform this methodology into a more scalable and reliable tool for educational research, enabling broader application across diverse educational contexts while maintaining analytical rigor.
-Acknowledgement
-We thank the editor, associate editor, and reviewers for their constructive comments.
-This work was partially supported by the National Research Foundation of Korea [grant number RS-2023-00217705 and RS-2024-00333701; Basic Science Research Program awarded to IHJ] and the ICAN (ICT Challenge and Advanced Network of HRD) support program [grant number RS-2023-00259934], supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation).
+Examining these teacher-student interactions would provide valuable insight into how instructional practices shape educational outcomes, particularly in the development of critical thinking skills and collaborative learning.
+Such an analysis could inform the optimization of teaching strategies and contribute to increased classroom effectiveness.
+An important avenue for future research involves exploring temporal variations in classroom interaction networks.
+Time-varying network approaches could effectively capture the evolution of interaction patterns throughout the instructional period, revealing how classroom dynamics change and develop over time.
+Such temporal analysis would illuminate whether the mediating effects of classroom interactions vary across different phases of the learning process, potentially yielding more nuanced insights into the dynamic nature of peer collaboration in the classroom and its impact on academic outcomes.
+Our current analysis faces methodological constraints that limit causal interpretation.
+Specifically, our mediation model incorporates only a limited set of confounding variables, relying primarily on pre-test performance as the primary confounding adjustment.
+This methodological limitation likely violates the “no omitted confounders” assumption essential for robust causal mediation interpretation.
+The collection of more comprehensive demographic, socioeconomic, and contextual variables would enable stronger causal inference and substantially improve the validity of our mediation findings.
+In addition, our current approach employs a two-stage estimation procedure due to inherent identifiability constraints within AMEN model specifications.
+Specifically, the rotation and reflection non-identifiability properties that characterize latent space models prevent reliable joint estimation of network formation and mediation parameters within a single unified framework.
+This methodological constraint introduces potential uncertainty propagation issues that may affect the precision of our mediation effect estimates.
+Future research should prioritize methodological enhancements that improve both the precision and generalizability of automated interaction analysis frameworks.
+While prompt engineering has demonstrated considerable promise, advancing beyond current approaches through model fine-tuning with high-quality labeled datasets would significantly enhance classification reliability in complex educational contexts.
+This advancement requires expanded human annotation efforts across diverse educational settings, systematically integrating expert pedagogical knowledge and carefully curated datasets into the model development process.
+Such comprehensive approaches would strengthen methodological rigor while preserving the efficiency advantages that make automated analysis attractive for large-scale educational research applications.
+A particularly critical advancement involves developing automated prompt generation methodologies that reduce human dependency in the analytical pipeline.
+Current approaches rely heavily on human judgment for prompt creation, optimization, and selection, introducing potential variability and bias that can compromise analytical consistency across different research contexts.
+Automating these prompt development processes would enhance both operational efficiency and methodological consistency, transforming LLM-based classification into a more scalable and standardized tool for educational research.
+Acknowledgment
+The authors thank the editor, associate editor, and reviewers for their constructive comments.
+The authors also thank Dr.
+Webb for providing the data and offering valuable insight for this study.
+This work was partially supported by the National Research Foundation of Korea [grant number NRF-2021S1A3A2A03088949, RS-2023-00217705, RS-2024-00333701; Basic Science Research Program awarded to IHJ], [grant number NRF-2025S1A5C3A0200632411; awarded to IHJ and MJ], International Joint Research Grant by Yonsei Graduate School [awarded to GK], and the ICAN (ICT Challenge and Advanced Network of HRD) support program [grant number RS-2023-00259934], supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation).
 Correspondence should be addressed to Ick Hoon Jin, Department of Applied Statistics, Department of Statistics and Data Science, Yonsei University, Seoul, Republic of Korea.